### v1

# Tensor-generated matrices and tensor H-eigenvalues distribution

Abstract
In this work, we introduce the concept of tensor-generated matrices, which transform an $m$-order $n$-dimensional tensor into an $n$-dimensional square matrix by grouping corresponding elements. We demonstrate that if the tensor-generated matrix from tensor $\mathcal{A}$ is an $H$-matrix, then $\mathcal{A}$ must be an $H$-tensor. While classical eigenvalue distributions for matrices are well-established, they do not directly apply to tensor eigenvalues, such as Brauer’s Ovals of Cassini sets, Ostrowski sets, and $S$-type inclusion sets. To overcome this challenge, we aim to convert high-order tensor $H$-eigenvalue localization into tensor-generated matrix eigenvalue localization. By establishing connections, we successfully obtain higher-order tensor $H$-eigenvalue distributions based on the subclass of $H$-matrices and matrix eigenvalue distributions. This approach enables us to extend existing matrix eigenvalue localization sets to higher-order tensor eigenvalues, resulting in modified versions of Brauer’s Ovals of Cassini sets, Ostrowski sets, and $S$-type inclusion sets.
keywords:
$H$-tensor, $H$-matrix, $H$-eigenvalue, Inclusion sets
1 Introduction
Tensors, also known as multidimensional arrays, and their eigenvalues have become increasingly important in the fields of applied mathematics and computational mathematics, fostering the development of numerical multilinear algebra. Moreover, tensors find wide-ranging applications in various domains, including liquid crystals [1, 2, 3, 4], higher-order Markov chains [5], best-rank one approximation in data analysis [6, 7, 8], quantum information [9, 10, 11, 12], and the positive definiteness of even-order multivariate forms in automatic control [13, 14].
Let $\mathbb{C}(\mathbb{R})$ represent the set of all complex (real) numbers, and let $N={1,2,\cdots,n}$ denote a positive integer $n\geq 2$. We define a complex (real) tensor $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})$ of order $m$ and dimension $n$ as $\mathbb{C}^{[m\times n]}(\mathbb{R}^{[m\times n]})$, where the entries satisfy
$a_{i_{1}i_{2}\cdots i_{m}}\in\mathbb{C}(\mathbb{R}),$
with $i_{j}=1,2,\cdots,n$ for $j=1,2,\cdots,m$. Notably, a vector corresponds to a tensor of order 1, and a matrix corresponds to a tensor of order 2. For a tensor $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{R}^{[m\times n]}$, $\mathcal{A}$ is considered nonnegative if every entry satisfies $a_{i_{1}i_{2}\cdots i_{m}}\geq 0$. A real tensor $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})$ is called symmetric if
$a_{i_{1}i_{2}\cdots i_{m}}=a_{\pi(i_{1}i_{2}\cdots i_{m})},\forall\pi\in\Pi_{m},$
where $\Pi_{m}$ represents the permutation group of $m$ indices. Furthermore, we refer to a real tensor of order $m$ and dimension $n$ as a unit tensor, denoted by $\mathcal{I}$, if its entries are given by $\delta_{i_{1}i_{2}\cdots i_{m}}$ for $i_{1},\cdots,i_{m}\in N$, where
$\delta_{i_{1}i_{2}\cdots i_{m}}=\begin{cases}1,&{\rm{if}}~{}i_{1}=\cdots=i_{m},\\ 0,&\rm{otherwise}.\end{cases}$
In 2005, Qi [15] and Lim [16] independently introduced the concept of eigenvalues for higher-order tensors. Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})$ be a complex tensor of order $m$ and dimension $n$, and let $x=(x_{1},x_{2},\cdots,x_{n})^{\top}$ be an $n$-dimensional vector. The expression $\mathcal{A}x^{m-1}$ represents an $n$-dimensional vector in $\mathbb{C}^{n}$, where its $i$-th component is defined as
$(\mathcal{A}x^{m-1})_{i}=\sum\limits_{i_{2},i_{3},\cdots,i_{m}=1}^{n}a_{ii_{2}\cdots i_{m}}x_{i_{2}}\cdots x_{i_{m}}.$
Furthermore, if there exist a complex number $\lambda$ and a nonzero complex vector $x=(x_{1},x_{2},\cdots,x_{m})^{\top}$ satisfying
$\mathcal{A}x^{m-1}=\lambda x^{[m-1]},$ (1)
then $\lambda$ is referred to as an $N$-eigenvalue of $\mathcal{A}$, and $x$ is an eigenvalue of $\mathcal{A}$ associated with $\lambda$, where
$x^{[m-1]}=(x_{1}^{m-1},x_{2}^{m-1},\cdots,x_{n}^{m-1})^{\top}.$
If both $\lambda$ and $x$ are real, then $\lambda$ is called an $H$-eigenvalue of $\mathcal{A}$, and $x$ is an $H$-eigenvector of $\mathcal{A}$ associated with $\lambda$ [15, 16].
We denote the set of all $H$-eigenvalues of $\mathcal{A}$ as $\sigma(\mathcal{A})$, which is referred to as the $H$-spectrum. The $H$-spectral radius of $\mathcal{A}$ is defined as:
$\rho(\mathcal{A})=\max{|\lambda|;\lambda\in\sigma(\mathcal{A})}.$
Consider an $m$-degree homogeneous polynomial form of $n$ variables, denoted as $f(x)$, where $x\in\mathbb{R}^{n}$ and $\mathcal{A}\in\mathbb{R}^{[m\times n]}$ in equation (2):
$f(x)=\mathcal{A}x^{m}=\sum\limits_{i_{1},i_{2},\cdots,i_{m}=1}^{n}a_{i_{1}i_{2}\cdots i_{m}}x_{i_{1}}x_{i_{2}}\cdots x_{i_{m}}.$ (2)
When $m$ is even, $f(x)$ is referred to as positive definite if
$f(x)>0~{}~{}~{}~{}\forall x\in\mathbb{R}^{n},x\neq 0.$
The positive definiteness of a multivariate polynomial $f(x)$ plays a significant role in studying the stability of nonlinear autonomous systems using Lyapunov’s direct method in automatic control [13]. Unfortunately, for $n\geq 3$ and $m\geq 4$, determining the positive definiteness of $f(x)$ becomes an NP-hard problem in mathematics. However, Qi [15] observed that $f(x)$ defined by (2) is positive definite if and only if the real symmetric tensor $\mathcal{A}$ is positive definite. Moreover, the tensor $\mathcal{A}$ is positive definite if and only if all $H$-eigenvalues of $\mathcal{A}$ are positive. Recently, Zhang [17] introduced $M$-tensors, which are extensions of $M$-matrices. The spectral theory of $M$-tensors provides a new method to certify the positive definiteness of multivariate forms. Additionally, Li et al. [18] proposed criteria for identifying the positive definiteness of even-order real symmetric tensors by considering $H$-tensors.
In addition, Azimzadeh and Bayraktar [19] have made notable advancements by extending the concept of weakly chained diagonal dominance from matrices to tensors. Their work demonstrates that a weakly chained diagonally dominant tensor with nonnegative diagonals can be classified as an $M$-tensor. Moreover, they have identified the significant theoretical implications of $M(H)$-tensors in solving high-order Bellman equations, which extend the classical Bellman equations to the tensor setting.
For instance, let us consider the following nonlinear problem: Find $u\in\mathbb{R}^{n}$ such that
$\min\limits_{P\in\mathcal{P}}{\mathcal{A}(P)u^{m-1}-b(P)}=0,$ (3)
where $\mathcal{A}(P)$ is an $m$-order and $n$-dimensional real tensor, $b(P)$ is a real vector, $\mathcal{P}=\mathcal{P}_{1}\times\cdots\times\mathcal{P}_{n}$ is a nonempty compact set, and the minimum is taken with respect to the coordinatewise order on $\mathbb{R}^{n}$. Here, each $\mathcal{P}=\mathcal{P}_{1}\times\cdots\times\mathcal{P}_{n}$ in $\mathcal{P}$ is an $n$-tuple with $\mathcal{P}_{i}$ belonging to $\mathcal{P}_{i}$. When the tensor $\mathcal{A}(P)$ is an $M$-tensor or a weakly chained diagonally dominant tensor, and subject to the conditions that $b(P)$ is a positive vector for each $P$ in $\mathcal{P}$ and $\mathcal{P}$ is finite, a positive solution $u$ exists for equation (3).
A tensor can be viewed as a complex nonlinear operator, and its product operation (including tensor-matrix and tensor-vector products) involves multiple indices, making it significantly more complex than matrix multiplication. With this motivation, our objective is to classify the elements of $m$th order $n$-dimensional tensors into square matrices of dimension $n$.
$H$-tensors play a vital role in scientific computation and engineering mathematics, such as in nonlinear control and polynomial equations. However, many existing methods for determining $H$-tensors are derived from generalizations of discriminant methods for $H$-matrices. These methods heavily rely on tensor-matrix and tensor-vector products, which greatly increase the complexity of the determination process.
To overcome this challenge, we propose transforming the problem of determining $H$-tensors into a discriminant problem for $H$-matrices. This transformation reduces the complexity of operations and determination. Moreover, by leveraging the direct use of the matrix generated by the tensor to determine $H$-tensors, almost all existing $H$-matrix determination methods can be extended to the determination of $H$-tensors. This approach will further advance and enrich the development of $H$-tensor theory.
Unlike the eigenvalue problem for matrices, computing the eigenvalues of higher-order tensors is known to be NP-hard. Nevertheless, several algorithms have been developed to calculate one or more eigenvalues of tensors, as documented in [20, 21, 22, 23, 24]. Unfortunately, these methods do not perform well for larger-sized tensors.
[…]
Theorem 1.
[15, Theorem 6] Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$, $n\geq 2$. Then
$\sigma(\mathcal{A})\subseteq\Gamma(\mathcal{A})=\bigcup_{i\in N}\Gamma_{i}(\mathcal{A}),$
where $\Gamma_{i}(\mathcal{A})=\{z\in\mathbb{C}:|z-a_{i\cdots i}|\leq r_{i}(\mathcal{A})\},\quad r_{i}(\mathcal{A})=\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|.$
[…]
Motivated by this, our work aims to transform the problem of higher-order tensor eigenvalue localization into the problem of tensor-generated matrix eigenvalue localization. By establishing the connection between subclasses of $H$-matrices and matrix eigenvalue distribution, we can determine the distribution of eigenvalues for higher-order tensors. In other words, we only need to locate the eigenvalues of the tensor-generated matrix to determine the positions of the original tensor eigenvalues. This approach allows us to extend the majority of matrix eigenvalue localization sets to higher-order tensor eigenvalues and present representative tensor eigenvalue inclusion sets, such as the classical Brauer’s Ovals of Cassini sets, Ostrowski sets, and $S$-type inclusion sets.
This paper focuses on exploring the relationships between an $m$-th order $n$-dimensional tensor and an $n$-dimensional square matrix. Specifically, we establish associations by classifying tensor elements with the same index and linking them to corresponding elements in the matrix. Based on these established relations, we introduce the concept of the tensor-generated matrix and demonstrate that if the tensor generated-matrix is an $H$-matrix, it implies that the original tensor must be an $H$-tensor.
Furthermore, we investigate several analogous properties shared by the original tensor and the tensor-generated matrix, such as weak irreducibility and weakly chained diagonal dominance. Additionally, we propose sufficient conditions for determining the positive definiteness of multivariate polynomials of even order. Leveraging the connections established between tensors and tensor-generated matrices, we extend classical matrix eigenvalue inclusion sets, which are typically challenging to extend to higher-order tensors, to higher-order tensor $H$-eigenvalues. This extension is facilitated through the connection between the $H$-matrix subclass and matrix eigenvalue distribution. Consequently, we propose representative tensor eigenvalue inclusion sets, including modified Brauer’s Ovals of Cassini sets, modified Ostrowski sets, and modified $S$-type inclusion sets of tensor $H$-eigenvalues.
The paper is organized as follows: Section 2 provides the necessary preliminaries. In Section 3, we introduce the concept of the tensor-generated matrix and establish the relationship between an $H$-matrix tensor-generated matrix and an $H$-tensor. We also explore similar properties shared by the original tensor and the tensor-generated matrix, such as weak irreducibility and weakly chained diagonal dominance. Moreover, we propose several sufficient conditions for determining the positive definiteness of multivariate polynomials of even order. In Section 4, we transform the problem of high-order tensor eigenvalue localization into the problem of tensor-generated matrix eigenvalue localization based on the established connections. We present representative tensor eigenvalue inclusion sets utilizing the connection between the $H$-matrix subclass and matrix eigenvalue distribution, including modified Brauer’s Ovals of Cassini sets, modified Ostrowski sets, and modified $S$-type inclusion sets. Finally, in Section 5, we provide numerical examples to illustrate our conclusions.
2 Preliminaries
In this section, we begin by introducing key definitions and established results regarding $H$-eigenvalues and $H$-tensors. It is important to note the consistent notation employed throughout this paper: lowercase letters, such as $x$, $y$, etc., represent vectors; italic capital letters, such as $A$, $B$, etc., represent matrices; and calligraphic capital letters, such as $\mathcal{A}$, $\mathcal{B}$, etc., denote tensors.
Definition 1.
[18, Definition 2] Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$. $\mathcal{A}$ is called an $H$-tensor if there is an entrywise positive vector $x=(x_{1},x_{2},\cdots,x_{n})^{\top}\in R^{n}$ such that for all $i\in N$,
$|a_{i\cdots i}|x_{i}^{m-1}>\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{i_{2}}\cdots x_{i_{m}}.$ (4)
Definition 2.
[17, Definition 3.14] A tensor $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$ is called a (strictly) diagonally dominant tensor if for $i\in N$,
$|a_{i\cdots i}|\geq(>)r_{i}(\mathcal{A})=\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|.$ (5)
And we denote
[…]
Definition 3.
[27, Definition 2.1] Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$. Then $\mathcal{A}$ is called reducible, if there exists a nonempty proper index $I\subset N$ such that
$a_{i_{1}i_{2}\cdots i_{m}}=0,\quad\forall i_{1}\in I,\quad\forall i_{2},\cdots,i_{m}\notin I.$
If $\mathcal{A}$ is not reducible, then $\mathcal{A}$ is called irreducible.
Definition 4.
[28, Definition 2.2] Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$ and $R(|\mathcal{A}|)$, the representation of $\mathcal{A}$, be a $n\times n$ matrix whose $(i,j)$-th entry is given by
$R(|\mathcal{A}|)_{ij}=\sum\limits_{j\in\{i_{2}\cdots i_{m}\}}|a_{ii_{2}\cdots i_{m}}|.$
Tensor $\mathcal{A}$ is called weakly reducible if $R(|\mathcal{A}|)$ is a reducible matrix. Otherwise, tensor $\mathcal{A}$ is weakly irreducible
There is another equivalent about weakly irreducible tensor. An order $m$ dimension $n$ tensor $\mathcal{A}$ is called weakly reducible, if there exists a nonempty proper index subset $I\subset N$ such that
$a_{i_{1}i_{2}\cdots i_{m}}=0,\quad\forall i_{1}\in I,\quad\exists i_{j}\notin I,\quad j=2,\cdots,m.$
If $\mathcal{A}$ is not weakly reducible, then we call $\mathcal{A}$ weakly irreducible [28, 29].
In the context of matrices, it is worth noting that there is no distinction between irreducible and weakly irreducible matrices. However, while an irreducible tensor must also be weakly irreducible, the reverse is not necessarily true. Therefore, the results applicable to weakly irreducible tensors can also be extended to irreducible tensors.
Definition 5.
[30, Definition 1.1] Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$, $D=diag(d_{1},$ $d_{2},\cdots,d_{n})$. Denote
$\mathcal{B}=(b_{i_{1}i_{2}\cdots i_{m}})=\mathcal{A}D^{m-1},\quad b_{i_{1}i_{2}\cdots i_{m}}=a_{i_{1}i_{2}\cdots i_{m}}d_{i_{2}}d_{i_{3}}\cdots d_{i_{m}},\quad i_{1},i_{2},\cdots,i_{m}\in N,$
we call $\mathcal{B}$ as the product of the tensor $\mathcal{A}$ and the matrix $X$.
Lemma 1.
[31, Proposition 29] Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$. Then $\mathcal{A}$ is an $H$-tensor if and only if exists a positive diagonal matrix $D$ such that $\mathcal{A}D^{m-1}$ is strictly diagonally dominant tensor.
Definition 6.
[31, Definition 3 and Definition 4] Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$. $\mathcal{A}$ is $Z$-tensor if all of its off-diagonal entries are non-positive, which is equivalent to write $\mathcal{A}=s\mathcal{I}-\mathcal{B}$, where $s>0$ and $\mathcal{B}$ is a nonnegative tensor. $\mathcal{A}$ is called an $M$-tensor if there exist a nonnegative tensor $\mathcal{B}$ and a positive real number $s>\rho(\mathcal{B})$ such that
$\mathcal{A}=s\mathcal{I}-\mathcal{B}.$
Lemma 2.
[29, Theorem 3.4] Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$. If $\mathcal{A}$ is irreducible,
$|a_{i\cdots i}|\geq r_{i}(\mathcal{A}),\quad\forall i\in N,$
and strictly inequality holds for at least one $i$, then $\mathcal{A}$ is an $H$-tensor.
Next, we introduce the directed graph associated with a tensor.
Definition 7.
[19, Definition 13] Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$ be a tensor.
[…]
Lemma 3.
[…]
Definition 8.
[19, Definition 15] A tensor $\mathcal{A}$ is weakly chained diagonally dominant if all of the following are satisfied
(i) $\mathcal{A}$ is diagonally dominant.
(ii) $J(\mathcal{A})$ is nonempty.
(iii) For each $i^{1}\notin J(\mathcal{A})$, there exists a walk $i^{1}\rightarrow i^{2}\rightarrow\cdots\rightarrow i^{k}$ in graph $\mathcal{A}$ such that $i^{k}\in J(\mathcal{A})$.
Lemma 4.
[19, Theorem 1] Let $\mathcal{A}$ be a weakly chained diagonally dominant $Z$-tensor with nonnegative diagonals, then $\mathcal{A}$ is a $M$-tensor.
Lemma 5.
[32, Lemma 2] Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$. If $\mathcal{A}$ is a $H$-tensor, then $0\notin\sigma(\mathcal{A})$.
3 The tensor-generated matrices and $H$-tensors
As a generalization of $H$-matrices to higher-order tensors, $H$-tensors also exhibit the property of pseudo-diagonal dominance. However, unlike the matrix-matrix product operation, the tensor-matrix product operation involves $m$ indices, resulting in significantly increased complexity compared to the matrix-matrix product. In other words, unlike matrices, most properties of pseudo-diagonal dominance, such as double diagonal dominance, $\gamma$-diagonal dominance, and $S$-diagonal dominance, cannot be directly extended to tensor cases.
Therefore, our focus is on transforming the determination of $H$-tensors into the determination of the $H$-matrix of the tensor-generated matrix. This matrix represents the entries of an $m$-th order $n$-dimensional tensor in an $n$-dimensional square matrix. This approach not only extends almost all properties of pseudo-diagonal dominance from matrices to higher-order tensors but also significantly reduces the number of index operations required for the tensor-matrix product.
In this section, we introduce the concept of tensor-generated matrices, which convert an $m$-th order $n$-dimensional tensor into an $n$-dimensional square matrix by grouping tensor elements with the same index. We demonstrate that if the tensor-generated matrix $A$ is an $H$-matrix, then the tensor $\mathcal{A}$ must be an $H$-tensor. Additionally, we investigate similar properties shared by the original tensor and the tensor-generated matrix, such as weak irreducibility and weakly chained diagonal dominance. Moreover, we propose a series of sufficient conditions for ensuring the positive definiteness of multivariate polynomials of even order.
For convenience’s sake, we present the following denotation in the remainder of the paper. We denote
$s_{ij}(\mathcal{A})=\frac{1}{m-1}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=j,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|,\quad j=1,2,\cdots,n.$
It is noting that $m$ are greater than or equal to 2. When $m=2$, $s_{ij}(A)$ becomes the row sum of the matrix minus the diagonal elements.
Definition 9.
Let $\mathcal{A}\in\mathbb{C}^{[m\times n]}$ be an $m$th-order $n$-dimensional tensor. We construct a square matrix $A=(a_{ij})\in\mathbb{C}^{n,n}$ with $n$ dimensions, defined as follows:
$a_{ij}=\begin{cases}|a_{ii\cdots i}|-s_{ii}(\mathcal{A}),&i=j,\\ s_{ij}(\mathcal{A}),&i\neq j.\end{cases}$ (6)
We refer to matrix $A\in\mathbb{C}^{n,n}$ as the tensor-generated matrix of tensor $\mathcal{A}$.
In this context, we represent the elements of an $m$th-order $n$-dimensional tensor $\mathcal{A}$ as a square matrix $A$ with $n$ dimensions by classifying the $n^{m}$ entries of $\mathcal{A}$. Next, we show that if the tensor-generated matrix $A\in\mathbb{C}^{n,n}$ is an $H$-matrix, then the tensor $\mathcal{A}$ is also an $H$-tensor.
Theorem 2.
Let $A\in\mathbb{C}^{n,n}$ be a tensor-generated matrix by tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. If matrix $A$ is an $H$-matrix, then the tensor $\mathcal{A}$ is an $H$-tensor.
Proof.
As we know, if the tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$ is an $H$-tensor, the following inequality (4) holds:
$|a_{i\cdots i}|x_{i}^{m-1}>\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{i_{2}}\cdots x_{i_{m}}.$
Using the inequality of arithmetic and geometric means, we can express the right-hand side of the above inequality as follows:
$\begin{split}&\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{i_{2}}\cdots x_{i_{m}}\\ =&\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|^{\frac{1}{m-1}}x_{i_{2}}\cdots|a_{ii_{2}\cdots i_{m}}|^{\frac{1}{m-1}}x_{i_{m}}\\ \leq&\frac{1}{m-1}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}(|a_{ii_{2}\cdots i_{m}}|x_{i_{2}}^{m-1}+\cdots+|a_{ii_{2}\cdots i_{m}}|x_{i_{m}}^{m-1})\\ =&\frac{1}{m-1}(\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{i_{2}}^{m-1}+\cdots+\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{i_{m}}^{m-1})\\ =&\frac{1}{m-1}\sum\limits_{t=2}^{m}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x^{m-1}_{i_{t}}.\end{split}$ (7)
It is note worthy that the following $m-1$ equalities hold for $p=2,\cdots,m$
$\begin{split}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}&|a_{ii_{2}\cdots i_{m}}|x_{i_{p}}^{m-1}=\sum\limits_{\begin{subarray}{c}i_{k}\in N,i_{p}=1,k=2\cdots m,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{1}^{m-1}\\ &+\cdots+\sum\limits_{\begin{subarray}{c}i_{k}\in N,i_{p}=n,k=2\cdots m,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{n}^{m-1},\end{split}$ (8)
If take $p=2$, then $m=2$ becomes a matrix, that is, $i_{3}\cdots i_{m}$ are vanished. In this situation, there is
$\sum\limits_{\delta_{ii_{2}}=0}|a_{ii_{2}}|x_{i_{2}}=\sum\limits_{\delta_{ii_{2}}=0}|a_{i1}|x_{1}+\cdots+\sum\limits_{\delta_{ii_{2}}=0}|a_{in}|x_{n}.$
We sum up the $q$-th term of all the above equalities for $q=1,\cdots,n$,
$\begin{split}\sum\limits_{\begin{subarray}{c}i_{k}\in N,i_{2}=q,k=2\cdots m,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}&|a_{ii_{2}\cdots i_{m}}|x_{q}^{m-1}+\cdots+\sum\limits_{\begin{subarray}{c}i_{k}\in N,i_{m}=q,k=2\cdots m,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{q}^{m-1}\\ =&\sum\limits_{k=2}^{n}\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=q,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{q}^{m-1}\end{split}$
Moreover, it follows that
$\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}\sum\limits_{t=2}^{m}|a_{ii_{2}\cdots i_{m}}|x^{m-1}_{i_{t}}=\sum\limits_{q=1}^{n}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=q,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{q}^{m-1}.$ (9)
Based on (7) and (9), there are
$\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{i_{2}}\cdots x_{i_{m}}\leq\frac{1}{m-1}\sum\limits_{t=2}^{m}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x^{m-1}_{i_{t}}=\sum\limits_{j=1}^{n}s_{ij}(\mathcal{A})x_{j}^{m-1}.$ (10)
In addition, the matrix $B=(b_{ij})=AD\in\mathbb{C}^{n,n}$ is defined as follows;
$b_{ij}=\begin{cases}(|a_{ii\cdots i}|-s_{ii}(\mathcal{A}))x_{i}^{m-1},&i=j,\\ s_{ij}(\mathcal{A})x_{j}^{m-1},&i\neq j,\end{cases}$ (11)
where $D=diag(x^{m-1}_{1},x^{m-1}_{2},\cdots,x^{m-1}_{n})$.
Since the tensor-generated matrix $A$ is an $H$-matrix, there exists a positive diagonal matrix
$D=\text{diag}(x^{m-1}{1},x^{m-1}{2},\cdots,x^{m-1}_{n})$ such that $B=AD$ is a strictly diagonally dominant matrix. In other words, we have:
$(|a_{ii\cdots i}|-s_{ii}(\mathcal{A}))x_{i}^{m-1}>\sum\limits_{j\neq i}^{n}s_{ij}(\mathcal{A})x_{j}^{m-1}.$
Then, the following inequalities are established by virtue of (10).
$\begin{split}|a_{ii\cdots i}|x_{i}^{m-1}>&\sum\limits_{j=1}^{n}s_{ij}(\mathcal{A})x_{j}^{m-1}\\ =&\frac{1}{m-1}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}\sum\limits_{t=2}^{m}|a_{ii_{2}\cdots i_{m}}|x^{m-1}_{i_{t}}\\ \geq&\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{i_{2}}\cdots x_{i_{m}}.\end{split}$
Therefore, according to Definition 1, tensor $\mathcal{A}$ is an $H$-tensor. The proof is completed. $\Box$
It is worth noting that when $x_{i_{2}}=\cdots=x_{i_{m}}=1$, that is, the equal in (7) holds, there are
$\begin{split}r_{i}(\mathcal{A})=&\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|=\frac{1}{m-1}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}\sum\limits_{t=2}^{m}|a_{ii_{2}\cdots i_{m}}|\\ =&\frac{1}{m-1}\sum\limits_{j=1}^{n}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=j,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|\\ =&s_{ii}(\mathcal{A})+\sum\limits_{j=1,j\neq i}^{n}s_{ij}(\mathcal{A}).\end{split}$ (12)
Clearly, based on the construction of the tensor-generated matrix $A\in\mathbb{C}^{n,n}$ from the tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$, we can draw the following conclusion.
Corollary 1.
Let $A\in\mathbb{C}^{n,n}$ be a tensor-generated matrix by tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. If the tensor-generated matrix $A$ is (strictly) diagonally dominant matrix, then tensor $\mathcal{A}$ is (strictly) diagonally dominant tensor.
Remark 1.
If there exists some $i_{0}\in N$ such that $|a_{i_{0}i_{0}\cdots i_{0}}|-s_{i_{0}i_{0}}(\mathcal{A})\leq 0$, it is evident that $\mathcal{A}$ is not a diagonally dominant tensor. Therefore, without special explanation, we always discuss diagonal dominance of tensors in the case of $|a_{ii\cdots i}|>s_{ii}(\mathcal{A})$ throughout this paper.
Remark 2.
According to equality (12) , the tensor with diagonally dominant in Definition 2 is consistent in the case of $|a_{ii\cdots i}|>s_{ii}(\mathcal{A})$.
For a $n$ dimension matrix $A=(a_{ij})\in\mathbb{C}^{n,n}$, we denote
$P_{i}(A)=\sum\limits_{j\neq i}|a_{ij}|,\quad Q_{i}(A)=\sum\limits_{j\neq i}|a_{ji}|.$
Therefore, for an $m$-order $n$-dimensional tensor $\mathcal{A}\in\mathbb{C}^{m\times n}$, based on the definition of the tensor-generated matrix $A=(a_{ij})\in\mathbb{C}^{n\times n}$, we can obtain $P_{i}(A)$ and $Q_{i}(A)$ as follows:
$P_{i}(A)=\sum\limits_{j=1,j\neq i}^{n}s_{ij}(\mathcal{A})=\frac{1}{m-1}\sum\limits_{j\neq i}^{n}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=j,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|,$
$Q_{i}(A)=\sum\limits_{i=1,i\neq j}^{n}s_{ij}(\mathcal{A})=\frac{1}{m-1}\sum\limits_{i\neq j}^{n}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=j,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|.$
It is well known that there are several forms of diagonal dominance for matrices. Here, we present examples of the most fundamental diagonally dominant matrices.
A matrix $A=(a_{ij})\in\mathbb{C}^{n,n}$ is considered to be diagonally dominant ($D_{n}$) if, for every $i\in N$, the absolute value of the diagonal element $|a_{ii}|$ is greater than the sum of the absolute values of the off-diagonal elements in the $i$th row:
$|a_{ii}|\geq P_{i}(A).$ (13)
$A=(a_{ij})\in\mathbb{C}^{n,n}$ is said to be a doubly diagonally dominant matrix($DD_{n}$) if for all $i,j\in N,i\neq j$,
$|a_{ii}||a_{jj}|\geq\big{(}P_{i}(A)\big{)}\big{(}P_{j}(A)\big{)}.$ (14)
$A=(a_{ij})\in\mathbb{C}^{n,n}$ is said to be a $\gamma$-diagonally dominant matrix($D^{\gamma}_{n}$) if there exists $\gamma\in[0,1]$ such that
$|a_{ii}|\geq\gamma\big{(}P_{i}(A)\big{)}+(1-\gamma)\big{(}Q_{i}(A)\big{)},\quad\forall i\in N.$ (15)
And $A=(a_{ij})\in\mathbb{C}^{n,n}$ is said to be a product $\gamma$-diagonally dominant matrix($PD^{\gamma}_{n}$) if there exists $\gamma\in[0,1]$ such that
$|a_{ii}|\geq\big{(}P_{i}(A)\big{)}^{\gamma}\big{(}Q_{i}(A)\big{)}^{1-\gamma},\quad\forall i\in N.$ (16)
If all the inequalities in (13)-(16) are strict, $A$ is classified as a strictly diagonally dominant matrix ($SD_{n}$), strictly doubly diagonally dominant matrix ($SDD_{n}$), strictly $\gamma$-diagonally dominant matrix ($SD^{\gamma}{n}$), and strictly product $\gamma$-diagonally dominant matrix ($SPD^{\gamma}_{n}$), respectively. Moreover, $A$ is referred to as a generalized diagonally dominant matrix ($GSD_{n}$), generalized $\gamma$-diagonally dominant matrix ($GSD^{\gamma}_{n}$), and generalized product $\gamma$-diagonally dominant matrix ($GSPD^{\gamma}_{n}$) if there exists a positive diagonal matrix $D$ such that $AD$ is a strictly diagonally dominant matrix, strictly $\gamma$-diagonally dominant matrix, and strictly product $\gamma$-diagonally dominant matrix, respectively. It is well known that matrix $A$ is a nonsingular $H$-matrix if there exists a positive diagonal matrix $D=(x_{1},\cdots,x_{n})^{\top}$ ($x_{i}>0$) for $i\in N$ such that $AD$ a strictly diagonally dominant matrix($SD_{n}$) [33]. Therefore, we have the following lemma hold. In particular, based on classical matrix eigenvalue inclusion set-Ostrowski set [33], there is the following item (5) (see Theorem 12 for details).
Lemma 6.
[34, 33] Let $A\in\mathbb{C}^{n,n}$. If matrix $A$ meets one of the following conditions, then matrix $A$ is a $H$-matrix.
(1) $A$ is a strictly diagonally dominant matrix ($SD_{n}$).
(2) $A$ is strictly doubly diagonally dominant matrix ($SDD_{n}$).
(3) $A$ is a strictly $\gamma$-diagonally dominant matrix ($SD^{\gamma}_{n}$).
(4) $A$ is a strictly product $\gamma$-diagonally dominant matrix ($SPD^{\gamma}_{n}$).
(5) $A$ is a generalized diagonally dominant matrix ($GSD_{n}$), generalized $\gamma$-diagonally dominant matrix ($GSD^{\gamma}_{n}$), and generalized product $\gamma$-diagonally dominant matrix ($GSPD^{\gamma}_{n}$).
However, apart from diagonal dominance, the aforementioned properties cannot be directly extended to higher-order tensors as they are applicable to matrices only.
Nonetheless, we can establish a connection between these diagonally dominant properties and higher-order tensors using the tensor-generated matrix. This allows us to bridge the gap between assessing $H$-tensors and $H$-matrices, which contributes to the advancement of $H$-tensor theory.
Similarly, we can demonstrate that the following classical matrix result can be extended to higher-order tensors using the tensor-generated matrix and Lemma 6.
Theorem 3.
Consider a tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$ that generates a matrix $A\in\mathbb{C}^{n,n}$. If the tensor-generated matrix $A$ satisfies any of the conditions outlined in Lemma 6, it implies that the tensor $\mathcal{A}$ is an $H$-tensor.
It is important to emphasize that the extension of basic diagonally dominant properties to higher-order tensors through the tensor-generated matrix is possible. Additionally, we can extend nearly all the determination results of $H$-matrices to higher-order tensors. While a detailed explanation is beyond the scope of this discussion, the previously mentioned classic conclusions serve as representative examples.
Furthermore, we can establish that the tensor-generated matrix $A$ and the tensor $\mathcal{A}$ share the same weak irreducibility. This enables us to extend the conclusions regarding $H$-matrices in the case of irreducibility to higher-order tensors.
Theorem 4.
Let $A\in\mathbb{C}^{n,n}$ be a tensor-generated matrix by tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. Then the tensor $\mathcal{A}$ is weakly irreducible, if and only if the tensor-generated matrix $A$ is irreducible.
Proof.
Necessity: $\mathcal{A}$ is a weakly irreducible tensor. We suppose that the tensor-generated matrix $A$ is reducible. Then there is a nonempty proper index subset $I\subset N$ such that $a_{ij}=0,\forall i\in I,\forall j\notin I$, that is to say that we have
$a_{ij}=s_{ij}(\mathcal{A})=\frac{1}{m-1}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{k-1},i_{k+1},\cdots,i_{m}\in N,i_{k}=j,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|=0,j\neq i.$
Furthermore, we obtain $a_{ii_{2}\cdots i_{m}}=0,\forall i\in I,\exists i_{2}\cdots i_{m}\notin I$, equivalently, $\mathcal{A}$ is a weakly reducible tensor. This leads to a contradiction.
Sufficiency: The tensor-generated matrix $A$ is irreducible. We assume that the tensor $\mathcal{A}$ is weakly reducible. In other words, there exists a nonempty proper index subset $I\subset N$ such that $a_{ii_{2}\cdots i_{m}}=0,\forall i\in I,\exists i_{2}\cdots i_{m}\notin I$, which implies that
$s_{ij}(\mathcal{A})=\frac{1}{m-1}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{k-1},i_{k+1},\cdots,i_{m}\in N,i_{k}=j,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|=0,j\neq i.$
Equivalently, there exists a nonempty proper index subset $I\subset N$ such that $a_{ij}=s_{ij}(\mathcal{A})=0,\forall i\in I,\forall j\notin I$, which is in contradiction with the irreducibility of matrix $A$. The proof is completed. ∎
Corollary 2.
Let $A\in\mathbb{C}^{n,n}$ be a tensor-generated matrix by tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. if the tensor $\mathcal{A}$ is irreducible, then the tensor-generated matrix $A$ is irreducible.
Based on Corollary 1, Theorem 4, and Lemma 2, we can readily derive the following conclusions.
Theorem 5.
Let $A\in\mathbb{C}^{n,n}$ be a tensor-generated matrix by tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. If the tensor-generated matrix $A$ is irreducible diagonally dominant matrix and strictly inequality holds for at least one, then tensor $\mathcal{A}$ is weakly irreducible diagonally dominant tensor. Moreover, tensor $\mathcal{A}$ is $H$-tensor.
Proof.
If the tensor-generated matrix $A\in\mathbb{C}^{n,n}$ by tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$ is irreducible, then Theorem 4 shows that tensor $A$ is weakly irreducible. Since the tensor-generated matrix $A\in\mathbb{C}^{n,n}$ is diagonally dominant, according to Corollary 1, tensor $\mathcal{A}$ is diagonally dominant. Moreover, on the basis of Lemma 2, tensor $\mathcal{A}$ is $H$-tensor. The proof is completed. ∎
Remark 3.
Since tensor $\mathcal{A}$ is weakly irreducible if and only if tensor-generated matrix $A\in\mathbb{C}^{n,n}$ by tensor $\mathcal{A}$ is irreducible, then we obtain the same conclusion as (Theorem 3.4 in [29]), that is, a weakly irreducible diagonally dominant tensor is a $H$-tensor.
Corollary 3.
Let $A\in\mathbb{C}^{n,n}$ be a tensor-generated matrix by irreducible tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. If the tensor-generated matrix $A$ is diagonally dominant matrix and strictly inequality holds for at least one, then tensor $\mathcal{A}$ is $H$-tensor.
Theorem 4 reveals that both the tensor and its tensor-generated matrix exhibit identical weak irreducibility. This observation allows us to focus our attention on studying the properties of $H$-tensors within the context of weak irreducibility.
Lemma 7.
[34, 33] Let $A\in\mathbb{C}^{n,n}$. If matrix $A$ meets one of the following conditions, then matrix $A$ is a $H$-matrix.
(1) $A$ is a irreducible diagonally dominant matrix and strictly inequality holds for at least one.
(2) $A$ is a irreducible doubly diagonally dominant matrix and strictly inequality holds for at least one.
(3) $A$ is a irreducible $\gamma$-diagonally dominant matrix and strictly inequality holds for at least one.
(4) $A$ is a irreducible product $\gamma$-diagonally dominant matrix and strictly inequality holds for at least one.
Theorem 5 establishes that the extension from case (1) in Lemma 7 to higher-order tensors is possible. However, this extension encompasses more than just that specific case. In fact, all cases presented in Lemma 7 can be generalized to the tensor case through the use of the tensor-generated matrix.
In other words, we can derive the following conclusions based on Lemma 7 and Theorem 2. The proof process follows a similar approach, and we will refrain from repeating it.
Theorem 6.
Let $A\in\mathbb{C}^{n,n}$ be a tensor-generated matrix by tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. If matrix $A$ satisfies one of the conditions in Lemma 7, then the tensor $\mathcal{A}$ is an $H$-tensor.
It is important to note that in the case of weak irreducibility, not only can we extend the scenarios presented in Lemma 7 to higher-order tensors, but we can also extend the conclusions of most matrix irreducible cases to higher-order tensors.
Furthermore, we can illustrate that if tensor $\mathcal{A}$ is a weakly chained diagonally dominant tensor, then the tensor-generated matrix $A$ is diagonally dominant. Additionally, there exists a non-zero element chain from $i\notin J(\mathcal{A})$ to $j\in J(\mathcal{A})$ for $i,j\in N$ (where $i\neq j$).
Theorem 7.
Let $A\in\mathbb{C}^{n,n}$ be a tensor-generated matrix by tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. If the tensor-generated matrix $A$ is a diagonally dominant matrix with a non-zero element chain from $i\notin J(\mathcal{A})$ to $j\in J(\mathcal{A})$ for $i,j\in N$, then tensor $\mathcal{A}$ is weakly chained diagonally dominant tensor. Moreover, tensor $\mathcal{A}$ is $H$-tensor.
Proof.
On the basis of Corollary 1, we obtain (i) and (ii) of Definition 8 straightway. And Lemma 3 implies that if the tensor-generated matrix $A$ have a non-zero element chain from $i\notin J(A)$ to $j\in J(A)$ for $i,j\in N$, the $\mathcal{A}$ also exists a walk $i^{1}\rightarrow i^{2}\rightarrow\cdots\rightarrow i^{k}$ in graph $\mathcal{A}$ such that $i^{k}\in J(\mathcal{A})$, where $i^{1}=i,i^{k}=j$. Hence, tensor $\mathcal{A}$ is weakly chained diagonally dominant by Definition 8. Based on Lemma 4, tensor $\mathcal{A}$ is $H$-tensor. This completes the proof. ∎
In short, in the case of strict inequality, weak irreducibility or weak chain, we can transform the determining of $H$-tensor into the determining of $H$-matrix. With this advantage, we can use the mature $H$-matrix theory to investigate more complex $H$-tensor problems, which greatly promotes the development of $H$-tensor theory. For example, we can use the spectral properties of $H$-tensors to derive sufficient conditions for positive definite tensors.
Theorem 8.
Let $\mathcal{A}\in\mathbb{C}^{[m\times n]}$ be an even-order real symmetric tensor with $a_{kk\cdots k}>0$ for all $k\in N$, and $A\in\mathbb{C}^{n,n}$ be a tensor-generated matrix by tensor $\mathcal{A}$. If matrix $A$ is a $H$-matrix, then the tensor $\mathcal{A}$ is positive definite.
In summary, when faced with strict inequality, weak irreducibility, or weak chains, we have the ability to convert the problem of determining an $H$-tensor into the problem of determining an $H$-matrix. This conversion provides us with the advantage of leveraging the well-established $H$-matrix theory to address more intricate problems involving $H$-tensors. As a result, the development of $H$-tensor theory is significantly advanced. For instance, we can utilize the spectral properties of $H$-tensors to derive sufficient conditions for positive definite tensors.
4 $H$-Eigenvalue inclusion sets for high order tensors
In this section, our focus is on the inclusion sets of $H$-Eigenvalues for higher-order tensors. It is worth noting that many classical eigenvalue localization results for matrices, such as Brauer’s Ovals of Cassini sets, Ostrowski sets, and $S$-type sets, are no longer applicable to tensor eigenvalues. To overcome this limitation, we aim to establish a connection between the localization of eigenvalues for higher-order tensors and the localization of eigenvalues for tensor-generated matrices, building upon the connection established in Section 3.
By leveraging this connection, we successfully determine the distribution of $H$-eigenvalues for higher-order tensors by relating it to the distribution of eigenvalues for the subclass of $H$-matrices. This approach allows us to extend the majority of matrix eigenvalue localization sets to higher-order tensor eigenvalues. In addition, we present some notable inclusion sets for tensor eigenvalues, such as the modified Brauer’s Ovals of Cassini sets, modified Ostrowski sets, and modified $S$-type inclusion sets.
[…]
Theorem 9.
Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$, $n\geq 2$. Then
${\sigma(\mathcal{A})}\subseteq{\mathcal{F}(\mathcal{A})}=\bigcup_{i\in N}{\mathcal{F}_{i}}(\mathcal{A}),$
where ${\mathcal{F}_{i}}(\mathcal{A})=\{z\in\mathbb{C}:[|z-a_{i\cdots i}|-s_{ii}(\mathcal{A})]\leq\sum\limits_{j=1,j\neq i}^{n}s_{ij}(\mathcal{A})\}$.
Proof.
For any $\lambda\in\sigma(\mathcal{A})$, let $\mathcal{B}=(b_{i_{1}i_{2}\cdots i_{m}})=\lambda\mathcal{I}-\mathcal{A}$. Then
$0\in\sigma(\mathcal{B}),\quad r_{i}(\mathcal{A})=r_{i}(\mathcal{B}),\quad s_{ij}(\mathcal{A})=s_{ij}(\mathcal{B}),\quad\forall i,j\in N.$
We assume $\lambda\notin\mathcal{F}(\mathcal{A})$, which indicate that there is $\lambda\notin\mathcal{F}_{i}(\mathcal{A}),\forall i\in N$. In other words, we obtain
$|\lambda-a_{i\cdots i}|-s_{ii}(\mathcal{A})>\sum\limits_{j=1,j\neq i}^{n}|s_{ij}(\mathcal{A})|,\quad\forall i\in N.$
Equivalently, we have
$|b_{i\cdots i}|-s_{ii}(\mathcal{B})>\sum\limits_{j=1,j\neq i}^{n}|s_{ij}(\mathcal{B})|,\quad\forall i\in N.$
According to equalities (12) and (13), $\mathcal{B}$ is a strictly diagonally dominant tensor. Moreover, there is $0\notin\sigma(\mathcal{B})$ by Lemma 5, which is in contradiction with the $0\in\sigma(\mathcal{B})$. The proof is completed. ∎
Remark 4.
[…]
Theorem 10 (Brauer’s Ovals of Cassini [33]).
Let $A=(a_{ij})\in\mathbb{C}^{n,n}$ be a $n\times n$ complex matrix, $n\geq 2$, and $\sigma(A)$ be the spectrum of $A$. Then
$\sigma(A)\subseteq\Theta(A)=\bigcup\limits_{\begin{subarray}{c}i,j\in N\\ j\neq i\end{subarray}}\Theta_{ij}(A),$
where $\Theta_{ij}(A)=\{z\in\mathbb{C}:|z-a_{ii}||z-a_{jj}|\leq r_{i}(A)r_{j}(A)\}$, $r_{i}(A)=\sum\limits_{k\neq i}|a_{ik}|$.
However, with the aid of the connections between tensors and tensor-generated matrices, along with the principles of double diagonal dominance and matrix eigenvalue distribution, we are able to obtain the following modified Brauer set for higher-order tensor $\mathcal{A}$.
Theorem 11.
Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$. Then
$\sigma(\mathcal{A})\subseteq\mathcal{D}(\mathcal{A})=\bigcup_{\begin{subarray}{c}i,j\in N\\ j\neq i\end{subarray}}\mathcal{D}_{i,j}(\mathcal{A}),$
where
$\mathcal{D}_{i,j}(\mathcal{A})=\{z\in\mathbb{C}:[|z-a_{ii\cdots i}|-s_{ii}(\mathcal{A})][|z-a_{jj\cdots j}|-s_{jj}(\mathcal{A})]\leq P_{i}(A)P_{j}(A)\}.$
Proof.
For any $\lambda\in\sigma(\mathcal{A})$, let $\mathcal{B}=(b_{i_{1}i_{2}\cdots i_{m}})=\lambda\mathcal{I}-\mathcal{A}$. Then
$0\in\sigma(\mathcal{B}),\quad r_{i}(\mathcal{A})=r_{i}(\mathcal{B}),\quad s_{ij}(\mathcal{A})=s_{ij}(\mathcal{B}),\quad\forall i,j\in N.$
We assume $\lambda\notin\mathcal{D}(\mathcal{A})$, which indicate that there is $\lambda\notin\mathcal{D}_{i,j}(\mathcal{A}),\forall i,j\in N$. In other words, we get
$[|\lambda-a_{ii\cdots i}|-s_{ii}(\mathcal{A})][|\lambda-a_{jj\cdots j}|-s_{jj}(\mathcal{A})]>P_{i}(A)P_{j}(A),\quad\forall i,j\in N,$
that is,
$[|b_{ii\cdots i}|-s_{ii}(\mathcal{B})][|b_{jj\cdots j}|-s_{jj}(\mathcal{B})]>P_{i}(B)P_{j}(B),\quad\forall i,j\in N.$
According to equalities (12) and (14), the tensor-generated matrix $B$ of tensor $\mathcal{B}$ is a strictly doubly diagonally dominant matrix, that is, $\mathcal{B}$ is a $H$-tensor. Moreover, there is $0\notin\sigma(\mathcal{B})$ by Lemma 5, which is in contradiction with the $0\in\sigma(\mathcal{B})$. The proof is completed. ∎
[…]
Theorem 12 (Ostrowski [33]).
Let $A=(a_{ij})\in\mathbb{C}^{n,n}$ be a $n\times n$ complex matrix, $n\geq 2$, $0\leq\gamma\leq 1$, and $\sigma(A)$ be the spectrum of $A$. Then
$\sigma(A)\subseteq\Omega(A)=\bigcup\limits_{i=1}^{n}\Omega_{ij}(A),$
where $\Theta_{ij}(A)=\{z\in\mathbb{C}:|z-a_{ii}|\leq(r_{i}(A))^{\gamma}(c_{i}(A))^{1-\gamma}\}$, $r_{i}(A)=\sum\limits_{j\neq i}|a_{ij}|$, $c_{i}(A)=\sum\limits_{j\neq i}|a_{ji}|$.
Similar to matrix eigenvalue Brauer sets, direct generalization of Ostrowski sets to higher-order tensors is not possible. However, by utilizing the connection between the tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$ and its corresponding tensor-generated matrix $A\in\mathbb{C}^{n,n}$, we can derive the modified Ostrowski sets for higher-order tensors $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. This is achieved by establishing a relationship between $\gamma$-diagonally dominant matrices and the distribution of matrix eigenvalues.
Theorem 13.
Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$, $n\geq 2$, $0\leq\gamma\leq 1$. Then
$\sigma(\mathcal{A})\subseteq\mathcal{O}(\mathcal{A})=\bigcup_{i\in N}\mathcal{O}_{i}(\mathcal{A}),$
where
$\mathcal{O}_{i}(\mathcal{A})=\{z\in\mathbb{C}:|z-|a_{i\cdots i}|+s_{ii}(\mathcal{A})|\leq[P_{i}(A)]^{\gamma}[Q_{i}(A)]^{1-\gamma}\}.$
Proof.
For any $\lambda\in\sigma(\mathcal{A})$, let $\mathcal{B}=(b_{i_{1}i_{2}\cdots i_{m}})=\lambda\mathcal{I}-\mathcal{A}$. Then
$0\in\sigma(\mathcal{B}),\quad r_{i}(\mathcal{A})=r_{i}(\mathcal{B}),\quad s_{ij}(\mathcal{A})=s_{ij}(\mathcal{B}),\quad\forall i,j\in N.$
We assume $\lambda\notin\mathcal{O}(\mathcal{A})$, which indicate that there is $\lambda\notin\mathcal{O}_{i}(\mathcal{A}),\forall i\in N$. In other words, we obtain
$[|\lambda-a_{i\cdots i}|-s_{ii}(\mathcal{A})]>[P_{i}(A)]^{\gamma}[Q_{i}(A)]^{1-\gamma},\quad\forall i\in N.$
That is to say
$[|b_{i\cdots i}|-s_{ii}(\mathcal{B})]>[P_{i}(B)]^{\gamma}[Q_{i}(B)]^{1-\gamma},\quad\forall i\in N.$
According to equalities (12) and (15), the tensor-generated matrix $B$ of tensor $\mathcal{B}$ is a strictly $\gamma$-diagonally dominant matrix, that is to say, $\mathcal{B}$ is a $H$-tensor. Moreover, there is $0\notin\sigma(\mathcal{B})$ by Lemma 5, which is in contradiction with the $0\in\sigma(\mathcal{B})$. The proof is completed. ∎
Lemma 8.
If $p,q>0$ and $0\leq\gamma\leq 1$, then
$p^{\gamma}q^{1-\gamma}\leq\gamma p+(1-\gamma)q.$
The equality holds if and only if $p=q$.
Immediately following Theorem 13 and Lemma 8, we derive the following Corollary.
Corollary 4.
Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$, $n\geq 2$, $0\leq\gamma\leq 1$. Then
$\sigma(\mathcal{A})\subseteq\mathcal{W}(\mathcal{A})=\bigcup_{i\in N}\mathcal{W}_{i}(\mathcal{A}),$
where
$\mathcal{W}_{i}(\mathcal{A})=\{z\in\mathbb{C}:[|z-a_{i\cdots i}|-s_{ii}(\mathcal{A})]\leq\gamma P_{i}(A)+(1-\gamma)Q_{i}(A)\}.$
Let set $S$ be any nonempty subset of $N=\{1,2,\cdots,n\}$, $n\geq 2$, where $\overline{S}=N\setminus S$ is the complement of the subset $S$ in $N$. For any given matrix $A=(a_{ij})\in\mathbb{C}^{n,n}$, each row sum $r_{i}(A)$ is divided into $S$ and $\overline{S}$ parts:
$\begin{cases}r_{i}(A)=\sum\limits_{j\in N\setminus\{i\}}|a_{ij}|=r_{i}^{S}(A)+r_{i}^{\overline{S}}(A),\\ r_{i}^{S}(A)=\sum\limits_{j\in S\setminus\{i\}}|a_{ij}|,\quad r_{i}^{\overline{S}}(A)=r_{i}(A)-r_{i}^{S}(A)=\sum\limits_{j\in\overline{S}\setminus\{i\}}|a_{ij}|,i\in N.\end{cases}$
Definition 10.
Let set $S$ be any nonempty subset of $N=\{1,2,\cdots,n\}$, $n\geq 2$, for any given matrix $A=(a_{ij})\in\mathbb{C}^{n,n}$, then the matrix $A$ is strictly $S$-diagonally dominant, if satisfied:
$\begin{cases}|a_{ii}|>r_{i}^{S}(A),\forall i\in S,\\ |a_{ii}|-r_{i}^{S}(A)\big{)}\cdot\big{(}|a_{jj}|-r_{j}^{\overline{S}}(A)\big{)}>r_{i}^{\overline{S}}(A)r_{j}^{S}(A),i\in S,j\in\overline{S}.\end{cases}$
The matrix eigenvalue $S$-type inclusion theorem is a prominent classical theory within the field of matrix eigenvalue inclusion theorems. This theorem directly follows from the matrix eigenvalue inclusion theorem based on $S$-diagonally dominant matrices.
Theorem 14.
[…]
$\Gamma_{i}^{S}(A)=\{z\in\mathbb{C}:|z-a_{ii}|\leq r_{i}^{S}(A)\},\forall i\in S,$
and the sets
$V_{i,j}^{S}(A)=\{z\in\mathbb{C}:\big{(}|z-a_{ii}|-r_{i}^{S}(A)\big{)}\cdot\big{(}|z-a_{jj}|-r_{j}^{\overline{S}}(A)\big{)}\leq r_{i}^{\overline{S}}(A)r_{j}^{S}(A)\},$
($\forall i\in S,\forall j\in\overline{S}$). Then,
$\sigma(A)\subseteq C^{S}(A)=\bigg{(}\bigcup_{i\in S}\Gamma_{i}^{S}(A)\bigg{)}\bigcup\bigg{(}\bigcup_{i\in S,j\in\overline{S}}V_{i,j}^{S}(A)\bigg{)}.$
$\begin{cases}P_{i}(\mathcal{A})=\sum\limits_{j\in N\setminus\{i\}}|a_{ij}|=r_{i}^{S}(\mathcal{A})+r_{i}^{\overline{S}}(\mathcal{A}),\\ r_{i}^{S}(\mathcal{A})=\sum\limits_{j\in S\setminus\{i\}}|s_{ij}|,\quad r_{i}^{\overline{S}}(\mathcal{A})=P_{i}(\mathcal{A})-r_{i}^{S}(\mathcal{A}),i\in N.\end{cases}$
Exploiting the established relationship between tensors and tensor-generated matrices in the previous section, we expand the classical matrix eigenvalue $S$-type inclusion theorem to encompass higher-order tensor $H$-eigenvalues. Consequently, we present a modified $S$-type inclusion theorem specifically tailored for tensor $H$-eigenvalues. This theorem enables us to effectively characterize the inclusion of $H$-eigenvalues within higher-order tensors by leveraging the connection between tensors and their tensor-generated matrices.
Theorem 15.
Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$ be a $m$ order $n$ dimension tensor and $S$ be any nonempty subset of $N=\{1,2,\cdots,n\}$, $n\geq 2$, where $\overline{S}=N\setminus S$ is the complement of the subset $S$ in $N$. Then
$\sigma(\mathcal{A})\subseteq\mathcal{C}^{S}(\mathcal{A})=\bigg{(}\bigcup_{i\in S}\Gamma_{i}^{S}(\mathcal{A})\bigg{)}\bigcup\bigg{(}\bigcup_{i\in S,j\in\overline{S}}\mathcal{C}_{i,j}^{S}(\mathcal{A})\bigg{)}$
where
$\Gamma_{i}^{S}(\mathcal{A})=\{z\in\mathbb{C}:||z-a_{i\cdots i}|-s_{ii}(\mathcal{A})|\leq r_{i}^{S}(\mathcal{A})\},\forall i\in S,$
$\begin{split}\mathcal{C}_{i,j}^{S}(\mathcal{A})=\{z\in\mathbb{C}:&[||z-a_{i\cdots i}|-s_{ii}(\mathcal{A})|-r_{i}^{S}(\mathcal{A})][||z-a_{j\cdots j}|-s_{jj}(\mathcal{A})|-r_{j}^{\overline{S}}(\mathcal{A})]\\ &\leq r_{i}^{\overline{S}}(\mathcal{A})r_{j}^{S}(\mathcal{A})\}.\end{split}$
Proof.
Let $\mathcal{B}=(b_{i_{1}i_{2}\cdots i_{m}})=\lambda\mathcal{I}-\mathcal{A}$, for any $\lambda\in\sigma(\mathcal{A})$, then
$0\in\sigma(\mathcal{B}),\quad r_{i}(\mathcal{A})=r_{i}(\mathcal{B}),\quad s_{ij}(\mathcal{A})=s_{ij}(\mathcal{B}),\quad\forall i,j\in N,$
and
$r_{i}^{S}(\mathcal{A})=r_{i}^{S}(\mathcal{B}),r_{i}^{\overline{S}}(\mathcal{A})=r_{i}^{\overline{S}}(\mathcal{B}),\quad\forall i\in N.$
We assume that $\lambda\notin\mathcal{C}^{S}(\mathcal{A})$, which shows that $\lambda\notin\bigg{(}\bigcup_{i\in S}\Gamma_{i}^{S}(\mathcal{A})\bigg{)}\bigcup\bigg{(}\bigcup_{i\in S,j\in\overline{S}}\mathcal{C}_{i,j}^{S}(\mathcal{A})\bigg{)},\forall i\in N$. In other words, we have
$\begin{cases}||\lambda-a_{i\cdots i}|-s_{ii}(\mathcal{A})|>r_{i}^{S}(\mathcal{A}),\\ [||\lambda-a_{i\cdots i}|-s_{ii}(\mathcal{A})|-r_{i}^{S}(\mathcal{A})][||\lambda-a_{j\cdots j}|-s_{jj}(\mathcal{A})|-r_{j}^{\overline{S}}(\mathcal{A})]>r_{i}^{\overline{S}}(\mathcal{A})r_{j}^{S}(\mathcal{A}),\end{cases}$
equivalently,
$\begin{cases}||b_{i\cdots i}|-s_{ii}(\mathcal{B})|>r_{i}^{S}(\mathcal{B}),\\ [||b_{i\cdots i}|-s_{ii}(\mathcal{B})|-r_{i}^{S}(\mathcal{B})][||b_{j\cdots j}|-s_{jj}(\mathcal{B})|-r_{j}^{\overline{S}}(\mathcal{B})]>r_{i}^{\overline{S}}(\mathcal{B})r_{j}^{S}(\mathcal{B}),\end{cases}$
From Definition 9 and Definition 10, the tensor-generated matrix $B$ of tensor $\mathcal{B}\in\mathbb{C}^{[m\times n]}$ is $S$-strictly diagonally dominant matrix. Hence, $\mathcal{B}$ is a $H$-tensor based on Theorem 2. Furthermore, it is result in $0\notin\sigma(\mathcal{B})$ by virtue of lemma 5, which is contradicted with $0\in\sigma(\mathcal{B})$. The proof is completed. ∎
When $S=\{i\}$ for $i\in N$, then Theorem 15 has the following case.
Corollary 5.
Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$. Then
$\sigma(\mathcal{A})\subseteq\mathcal{C}^{S^{i}}(\mathcal{A})=\bigcup_{i\in S,j\in\overline{S}}\mathcal{C}_{i,j}^{S^{i}}(\mathcal{A})$
where
$\mathcal{C}_{i,j}^{S^{i}}(\mathcal{A})=\{z\in\mathbb{C}:||z-a_{i\cdots i}|-s_{ii}(\mathcal{A})|\big{(}||z-a_{j\cdots j}|-s_{jj}(\mathcal{A})|-P_{j}(\mathcal{A})+s_{ji}(\mathcal{A})\big{)}\leq P_{i}(\mathcal{A})s_{ji}(\mathcal{A})\}.$
[…]
Theorem 4 reveals the consistent relationship between the irreducibility of the tensor-generated matrix and the weak irreducibility of the original tensor. Consequently, when dealing with irreducible tensor-generated matrices, our method is also suitable for Brualdi-type inclusion sets [38, 33]. Although we will omit the derivation process here, similar principles are applied.
To summarize, in Section 4, we transform the problem of tensor eigenvalue localization into matrix eigenvalue localization. This allows us to leverage the mature theory of matrix eigenvalue localization and apply it to the investigation of tensor eigenvalue localization.
5 Numerical examples
In this section, we provide numerical examples to demonstrate the validity and effectiveness of the tensor $H$-eigenvalue inclusion theorem derived in the previous section.
We begin by examining Example 1.1 from [35], which serves to illustrate that the concept of Brauer’s Ovals of Cassini sets, commonly associated with matrix eigenvalues, cannot be directly extended to an inclusion theorem for higher-order tensor $H$-eigenvalues.
Example 1.
Let $\mathcal{A}\in\mathbb{R}^{[4\times 2]}$, the non-zero elements are as follows
$a_{1111}=7,a_{1112}=a_{1121}=a_{1211}=a_{2111}=-2;$
$a_{2222}=6,a_{2221}=a_{2212}=a_{2122}=a_{1222}=-1;$
otherwise $a_{i_{1}i_{2}i_{3}i_{4}}=0$.
For a 4 order 2 dimension tensor, we can correspond the elements of this tensor to a $2\times 2$ tensor-generated matrix (6) by virtue of the connection between the tensor and the tensor-generated matrix. Hence, we obtian
$s_{11}=4,s_{12}=3,s_{21}=3,s_{22}=2,$
and
$A=\left(\begin{array}[]{cc}|7|-|4|&3\\ 3&|6|-|2|\\ \end{array}\right).$
From Theorem 11, we get
$\sigma(\mathcal{A})\subseteq\mathcal{\bar{O}}(\mathcal{A})=\{z\in\mathbb{C}:||z-7|-4|||z-6|-2|\leq 3\times 3\},$
and
$\sigma(\mathcal{A})\subseteq\mathcal{\bar{O}}(\mathcal{A})=\{z\in\mathbb{C}:0.4586\leq z\leq 12.8541\}.$
In fact, the eigenvalues of tensor $\mathcal{A}=(a_{i_{1}i_{2}i_{3}i_{4}})$ are
$\sigma(\mathcal{A})=\{12.7389,0.4725\}.$
It shows that the modified ovality theorem of higher order tensor eigenvalues is correct.
Example 2.
Let $\mathcal{A}\in\mathbb{R}^{[4\times 4]}$ be a 4 order 4 dimension real tensor whose elements are defined as follows:
$a_{1111}=10,\quad a_{2222}=8,\quad a_{3333}=7,\quad a_{4444}=5,$
$a_{1333}=a_{1444}=1,\quad a_{1211}=a_{1113}=a_{1141}=1,$
$a_{1332}=a_{1442}=a_{1232}=a_{1234}=a_{1321}=a_{1214}=1,$
$a_{2333}=a_{2444}=1,\quad a_{2112}=a_{2234}=a_{2113}=a_{2343}=a_{2123}=1,$
$a_{3222}=1,\quad a_{3111}=1,a_{3121}=a_{3434}=a_{3123}=1,$
$a_{4222}=1,\quad a_{4111}=1,a_{4121}=a_{4334}=1,$
otherwise $a_{i_{1}i_{2}i_{3}i_{4}}=0$.
We can correspond the elements of $\mathcal{A}\in\mathbb{R}^{[4\times 4]}$ to a $4\times 4$ tensor-generated matrix (6), then
$s_{11}=\frac{8}{3},s_{12}=\frac{8}{3},s_{13}=3,s_{14}=\frac{8}{3},s_{21}=\frac{5}{3},s_{22}=1,s_{23}=\frac{8}{3},s_{24}=\frac{5}{3},$
$s_{31}=2,s_{32}=\frac{5}{3},s_{33}=\frac{2}{3},s_{34}=\frac{2}{3},s_{41}=\frac{5}{3},s_{42}=\frac{4}{3},s_{43}=\frac{2}{3},s_{44}=\frac{1}{3},$
and
$A=\left(\begin{array}[]{cccc}10-\frac{8}{3}&\frac{8}{3}&3&\frac{8}{3}\\ \frac{5}{3}&8-1&\frac{8}{3}&\frac{5}{3}\\ 2&\frac{5}{3}&7-\frac{2}{3}&\frac{2}{3}\\ \frac{5}{3}&\frac{4}{3}&\frac{2}{3}&5-\frac{1}{3}\\ \end{array}\right).$
First, the results of some existing eigenvalue inclusion theorem are shown as following. From Theorem 6 in [15], one has
$\sigma(\mathcal{A})\subseteq\Gamma(\mathcal{A})=\{z\in\mathbb{C}:-1\leq z\leq 21\}.$
By virtue of Theorem 2.2 in [35], one has
$\sigma(\mathcal{A})\subseteq\mathcal{K}^{S}(\mathcal{A})=\{z\in\mathbb{C}:-0.7016\leq z\leq 20.3739\}.$
Based on Theorem 4 in [39], one has
$\sigma(\mathcal{A})\subseteq\Omega^{S}(\mathcal{A})=\{z\in\mathbb{C}:-0.4641\leq z\leq 19.7823\}.$
In the light of Theorem 11 in [32], one has
$\sigma(\mathcal{A})\subseteq\Phi(\mathcal{A})=\{z\in\mathbb{C}:-0.3723\leq z\leq 19.6847\}.$
Next, we will present our results of eigenvalue inclusion theorem. According to Theorem 9, we have
$\sigma(\mathcal{A})\subseteq\mathcal{F}(\mathcal{A})=\{z\in\mathbb{C}:-1\leq z\leq 21\}.$
It can be seen that the modified disk Theorem obtained from tensor-generated matrix is consistent with the result obtained from Theorem 6 in [15].
On the basis of Theorem 11, we obtain
$\sigma(\mathcal{A})\subseteq\mathcal{D}(\mathcal{A})=\{z\in\mathbb{C}:0.0936\leq z\leq 18.1382\}.$
From Theorem 13, we take $\gamma=0.5$, then
$\sigma(\mathcal{A})\subseteq\mathcal{O}(\mathcal{A})=\{z\in\mathbb{C}:0.3849\leq z\leq 19.3333\}.$
If we take $\gamma=0.04$, then
$\sigma(\mathcal{A})\subseteq\mathcal{O}(\mathcal{A})=\{z\in\mathbb{C}:-0.2717\leq z\leq 18.0961\}.$
On the grounds of Corollary 4, we take $\gamma=0.5$, then
$\sigma(\mathcal{A})\subseteq\mathcal{W}(\mathcal{A})=\{z\in\mathbb{C}:0.3333\leq z\leq 19.5000\}.$
If we take $\gamma=0.04$, then
$\sigma(\mathcal{A})\subseteq\mathcal{W}(\mathcal{A})=\{z\in\mathbb{C}:-0.2800\leq z\leq 18.1200\}.$
If we take $S=\{1,2\}$, we can easily get the following result by calculation.
$r_{1}^{S}(\mathcal{A})=\frac{8}{3},r_{1}^{\overline{S}}(\mathcal{A})=\frac{17}{3},r_{2}^{S}(\mathcal{A})=\frac{5}{3},r_{2}^{\overline{S}}(\mathcal{A})=\frac{13}{3},$
$r_{3}^{S}(\mathcal{A})=\frac{11}{3},r_{3}^{\overline{S}}(\mathcal{A})=\frac{2}{3},r_{4}^{S}(\mathcal{A})=3,r_{4}^{\overline{S}}(\mathcal{A})=\frac{2}{3}.$
From Theorem 15, it is follow as
$\sigma(\mathcal{A})\subseteq\mathcal{C}^{S}(\mathcal{A})=\{z\in\mathbb{C}:0.5811\leq z\leq 17.5803\}.$
Based on Corollary 5, one has
$\sigma(\mathcal{A})\subseteq\mathcal{C}^{S^{i}}(\mathcal{A})=\{z\in\mathbb{C}:-0.4741\leq z\leq 19.8130\}.$
In fact, we can compute all the $H$-eigenvalues of the tensor $\mathcal{A}$ as follow:
$\sigma(\mathcal{A})=\{4.4858,7.3107,9.7718,15.2641\}.$
This demonstrates that our eigenvalue localization set, consisting of Theorem 9, Theorem 11, Corollary 4, Theorem 15, and Corollary 5, successfully captures all the eigenvalues. Furthermore, the results obtained from our methods in the aforementioned examples reveal that the modified Brauer-type sets, modified Ostrowski sets, and modified $S$-type inclusion sets for tensor eigenvalues obtained from the tensor-generated matrix offer superior precision compared to existing sets for tensor eigenvalues.
Conclusions
In this study, we present a novel approach that focuses on the tensor-generated matrix of $H$-tensors and establishes the $H$-eigenvalue inclusion theorem for higher-order tensors. Firstly, we introduce the concept of the tensor-generated matrix, which converts an $m$-order $n$-dimensional tensor into an $n$-dimensional square matrix. We demonstrate that if the tensor-generated matrix derived from tensor $\mathcal{A}$ is an $H$-matrix, then tensor $\mathcal{A}$ must be an $H$-tensor. Furthermore, we investigate similar properties of the original tensor and the tensor-generated matrix, such as weak irreducibility and weakly chained diagonal dominance. Additionally, we propose a series of sufficient conditions for ensuring the positive definiteness of even-order multivariate polynomials.
Moreover, we establish connections between high-order tensor eigenvalue localization and tensor-generated matrix eigenvalue localization. By leveraging these connections, we present the distribution of higher-order tensor $H$-eigenvalues based on the subclass of $H$-matrices and matrix eigenvalue distribution. This approach allows us to extend the majority of matrix eigenvalue localization techniques to higher-order tensor eigenvalues. We introduce several representative sets for tensor eigenvalue inclusion, such as the modified Brauers Ovals of Cassini sets, modified Ostrowski sets, and modified S-type inclusion sets. Finally, we provide numerical examples to illustrate the correctness and superiority of our proposed results. Through these examples, we demonstrate the practical effectiveness and reliability of our approach.
Data availibility
All data generated or analysed during this study are included in this published article.
Declarations
Conflict of interest The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.