### v2

# The Role of Tensor-Generated Matrices in Analyzing Spin State Classicality and Tensor H-Eigenvalue Distributions

Abstract
Multipartite quantum scenarios are a significant and challenging resource in quantum information science. Tensors provide a powerful framework for representing multipartite quantum systems. In this work, we introduce the role of tensor-generated matrices that can broadly be defined as the relationships between an $m$-th order $n$-dimensional tensor and an $n$-dimensional square matrix. Through these established connections, we demonstrate that the classification of the tensor-generated matrix as an $H$-matrix implies the original tensor is also an $H$-tensor. We also explore various similar properties exhibited by both the original tensor and the tensor-generated matrix, including weak irreducibility, weakly chained diagonal dominance, and (strong) symmetry. These findings provide a method to transform intricate tensor problems into matrices in specific contexts, which is especially pertinent due to the NP-hard complexity of the majority of tensor problems. Subsequently, we explore the application of tensor-generated matrices in analyzing the classicality of spin states. Leveraging the tensor representation, we introduce classicality criteria for (strongly) symmetric spin-$j$ states, which potentially provide fresh perspectives on the study of multipartite quantum resources. Finally, we extend classical matrix eigenvalue inclusion sets to higher-order tensor $H$-eigenvalues, a task that is typically challenging for higher-order tensors. Consequently, we propose representative tensor $H$-eigenvalue inclusion sets, such as modified Brauer’s Ovals of Cassini sets, Ostrowski sets, and $S$-type inclusion sets.
keywords:
$H$-tensor, quantum information, spin coherent state, Classicality (Separability), $H$-eigenvalue
journal: SIAM Journal on Matrix Analysis and Applications
1 Introduction
Tensors, also known as multidimensional arrays, and their eigenvalues have become increasingly important in the fields of applied mathematics and computational mathematics, fostering the development of numerical multilinear algebra. Moreover, tensors find wide-ranging applications in various domains, including liquid crystals [1, 2, 3, 4], higher-order Markov chains [5], best-rank one approximation in data analysis [6, 7, 8], quantum information [9, 10, 11, 12, 13], and the positive definiteness of even-order multivariate forms in automatic control [14, 15].
With the progress in technology and the evolution of quantum platforms, many platforms are now capable of concurrently generating and manipulating multipartite qubits. It has become evident in recent years that high-dimensional systems offer advantages in specialized tasks such as quantum communication [16, 17] and quantum computing [18, 19]. Consequently, an increasing amount of multi-order high-dimensional tensor data will manifest in multipartite quantum systems.
Tensors provide a robust framework for representing complex quantum data structures [20, 21, 22, 23]. Analyzing the eigenvalues of these tensors uncovers crucial insights into quantum system properties and behaviors [9, 10]. This exploration is essential for understanding multipartite quantum states [24], entanglement phenomena, and optimizing quantum computational tasks [25]. By deciphering the properties of tensors [26, 27], we can reveal intricate quantum correlations, leading to advancements in quantum algorithms and communication protocols. This research not only enriches the theoretical foundations of quantum information science but also drives innovations in quantum technologies, enabling more efficient information processing and communication capabilities in various practical applications.
The coherent superposition of states [28], along with observable quantization, stands out as a key characteristic signifying the distinction of quantum mechanics from classical physics. Quantum coherence within many-body systems encapsulates the core nature of entanglement and serves as a crucial element for numerous physical phenomena in quantum optics [29, 30], quantum information [28], solid-state physics [31], and nanoscale thermodynamics [32]. Recently, there has been a growing interest in studying the existence and functional significance of quantum coherence [33, 34, 35]. Despite its fundamental significance, the establishment of a comprehensive theory regarding quantum coherence as a physical asset has only commenced in recent times.
Spin coherent states are the ones that come as close as possible to the ideal of a classical phase-space point, in the sense that their quantum fluctuations for the angular moment components are as small as allowed by Heisenbergs uncertainty relation [36], which is widely used in quantum optics [29, 30] and other fields [37]. The spin-$j$ states is classical [38, 39] if its density matrix can be decomposed as a weighted sum of angular momentum coherent states with positive weights, that is, the set of classical spin states are defined as the convex hull of spin coherent states. Quantum states that lie beyond the domain of classical spin states are regarded as nonclassical (truly quantum). The quantumness [40, 41] of a spin-$j$ state, used to quantify the separation between a quantum state and a classical one, can be gauged by its distance from the classical states. Notably, employing various distance metrics will yield different measures of quantumness, including the Hilbert-Schmidt distance [40], Bures distance [42], and trace distance [43].
Despite tensors’ superior descriptive capacity for multi-order high-dimensional data compared to matrices, a notable challenge arises from the prevalence of NP-hard problems in tensor analysis [44]. In contrast to matrix eigenvalue computations, determining eigenvalues and confirming positive semidefiniteness in higher-order tensors is acknowledged to be NP-hard. Despite ongoing efforts, various algorithms have emerged for calculating tensor eigenvalues, as detailed in [45, 46, 47, 48, 49]. Regrettably, these methods exhibit limitations in handling larger tensors.
[…]
In this paper, we introduce the role of tensor-generated matrices and their applications in analyzing spin state classicality and tensor $H$-Eigenvalue distributions. Firstly, the tensor-generated matrices can broadly be defined as the relationships between an $m$-th order $n$-dimensional tensor $\mathcal{A}$ and an $n$-dimensional square matrix $A$. Specifically, this classification allows for aligning the elements of tensor $\mathcal{A}$ with the corresponding elements of matrix $A$ of identical dimensions. Through these established relationships, we show that when the tensor-generated matrix is classified as an $H$-matrix, it necessarily infers that the original tensor is an $H$-tensor. We also investigate several analogous properties shared by the original tensor and the tensor-generated matrix, such as weak irreducibility, weakly chained diagonal dominance and (strong) symmetry. The findings will facilitate an extensive exploration of $H$-tensor theory by leveraging the well-established $H$-matrix theory. Additionally, they offer a methodology to convert complex tensor issues into matrices in specific scenarios, particularly relevant given the NP-hard nature of most tensor problems.
Subsequently, we investigate the use of tensor-generated matrices for analyzing the classicality of spin states. The preceding analysis establishes ample criteria for determining the semi-positive definiteness of even-order symmetric tensors. Drawing on the tensor representation of spin states and the tensor-generated matrix, we introduce specific classicality criteria for spin-$j$ states. Notably, an important observation arises: when the original tensor displays strong symmetry, the resulting tensor-generated matrix also exhibits symmetry. Consequently, we derive classicality criteria for spin-$j$ states characterized by strongly symmetric coefficient tensors.
Finally, by leveraging the established connections between tensors and tensor-generated matrices, we extend classical matrix eigenvalue inclusion sets to higher-order tensor $H$-eigenvalues, a task that is typically challenging for higher-order tensors. This extension is enabled by the relationship between the $H$-matrix subclass and matrix eigenvalue distribution. Consequently, we propose representative tensor eigenvalue inclusion sets, such as modified Brauer’s Ovals of Cassini sets, modified Ostrowski sets, and modified $S$-type inclusion sets of tensor $H$-eigenvalues.
The paper is organized as follows: Section 2 provides the necessary preliminaries for tensor and quantum information theory. In Section 3, we introduce the concept of the tensor-generated matrix and establish the relationship between an $H$-matrix tensor-generated matrix and an $H$-tensor. We also explore similar properties shared by the original tensor and the tensor-generated matrix, such as weak irreducibility and weakly chained diagonal dominance. In section 4, we explore the tensor-generated matrix application in spin states, and present some classical criteria for (strongly) symmetric spin-$j$ states. In Section 5, based on the established connections, we present representative tensor eigenvalue inclusion sets utilizing the connection between the $H$-matrix subclass and matrix eigenvalue distribution, including modified Brauer’s Ovals of Cassini sets, modified Ostrowski sets, and modified $S$-type inclusion sets.
2 Preliminaries
In this section, we begin by introducing key definitions and established results regarding $H$-tensors and quantum information. It is important to note the consistent notation employed throughout this paper: lowercase letters, such as $x$, $y$, etc., represent vectors; italic capital letters, such as $A$, $B$, etc., represent matrices; and calligraphic capital letters, such as $\mathcal{A}$, $\mathcal{B}$, etc., denote tensors.
2.1 Preliminaries for tensor
Let $\mathbb{C}(\mathbb{R})$ represent the set of all complex (real) numbers, and let $N={1,2,\cdots,n}$ denote a positive integer $n\geq 2$. We define a complex (real) tensor $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})$ of order $m$ and dimension $n$ as $\mathbb{C}^{[m\times n]}(\mathbb{R}^{[m\times n]})$, where the entries satisfy $a_{i_{1}i_{2}\cdots i_{m}}\in\mathbb{C}(\mathbb{R}),$ with $i_{j}=1,2,\cdots,n$ for $j=1,2,\cdots,m$. Notably, a vector corresponds to a tensor of order 1, and a matrix corresponds to a tensor of order 2. For a tensor $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{R}^{[m\times n]}$, $\mathcal{A}$ is considered nonnegative if every entry satisfies $a_{i_{1}i_{2}\cdots i_{m}}\geq 0$. A real tensor $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})$ is called symmetric if
$a_{i_{1}i_{2}\cdots i_{m}}=a_{\pi(i_{1}i_{2}\cdots i_{m})},\forall\pi\in\Pi_{m},$
where $\Pi_{m}$ represents the permutation group of $m$ indices. A unit tensor, denoted by $\mathcal{I}$, has all diagonal entries 1 and otherwise entries 0.
In 2005, Qi [52] and Lim [53] independently introduced the concept of eigenvalues for higher-order tensors. Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})$ be a complex tensor of order $m$ and dimension $n$, and let $x=(x_{1},x_{2},\cdots,x_{n})^{\top}$ be an $n$-dimensional vector. The expression $\mathcal{A}x^{m-1}$ represents an $n$-dimensional vector in $\mathbb{C}^{n}$, where its $i$-th component is defined as
$(\mathcal{A}x^{m-1})_{i}=\sum\limits_{i_{2},i_{3},\cdots,i_{m}=1}^{n}a_{ii_{2}\cdots i_{m}}x_{i_{2}}\cdots x_{i_{m}}.$
If there exist a complex number $\lambda$ and a nonzero complex vector $x=(x_{1},x_{2},\cdots,x_{m})^{\top}$ satisfying
$\mathcal{A}x^{m-1}=\lambda x^{[m-1]},$ (1)
then $\lambda$ is referred to as an $N$-eigenvalue of $\mathcal{A}$, and $x$ is an eigenvalue of $\mathcal{A}$ associated with $\lambda$, where $x^{[m-1]}=(x_{1}^{m-1},x_{2}^{m-1},\cdots,x_{n}^{m-1})^{\top}.$ If both $\lambda$ and $x$ are real, then $\lambda$ is called an $H$-eigenvalue of $\mathcal{A}$, and $x$ is an $H$-eigenvector of $\mathcal{A}$ associated with $\lambda$ [52, 53]. We denote the set of all $H$-eigenvalues of $\mathcal{A}$ as $\sigma(\mathcal{A})$, which is referred to as the $H$-spectrum. The $H$-spectral radius of $\mathcal{A}$ is defined as:
$\rho(\mathcal{A})=\max\{|\lambda|;\lambda\in\sigma(\mathcal{A})\}.$
Consider an $m$-degree homogeneous polynomial form of $n$ variables, denoted as $f(x)$, where $x\in\mathbb{R}^{n}$ and $\mathcal{A}\in\mathbb{R}^{[m\times n]}$ in equation (2):
$f(x)=\mathcal{A}x^{m}=\sum\limits_{i_{1},i_{2},\cdots,i_{m}=1}^{n}a_{i_{1}i_{2}\cdots i_{m}}x_{i_{1}}x_{i_{2}}\cdots x_{i_{m}}.$ (2)
When $m$ is even, $f(x)$ is referred to as positive semidefinite if
$f(x)\geq 0~{}~{}~{}~{}\forall x\in\mathbb{R}^{n},x\neq 0.$
In general, for $n\geq 3$ and $m\geq 4$, determining the positive definiteness of $f(x)$ becomes an NP-hard problem in mathematics. However, Qi [52] observed that $f(x)$ defined by (2) is positive definite if and only if the real symmetric tensor $\mathcal{A}$ is positive semidefinite.
The product of a tensor $\mathcal{A}$ and a diagonal matrix $D$ is defined as follows [54]: Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{m\times n}$ and $D=\text{diag}(d_{1},d_{2},\ldots,d_{n})$, one has
$\mathcal{B}=(b_{i_{1}i_{2}\cdots i_{m}})=\mathcal{A}D^{m-1},\quad b_{i_{1}i_{2}\cdots i_{m}}=a_{i_{1}i_{2}\cdots i_{m}}d_{i_{2}}d_{i_{3}}\cdots d_{i_{m}},\quad i_{1},i_{2},\cdots,i_{m}\in N.$
The resulting tensor $\mathcal{B}$ is termed as the product of tensor $\mathcal{A}$ and the matrix $D$.
The tensor $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$ is termed an $H$-tensor [55] if there exists an entrywise positive vector $x=(x_{1},x_{2},\ldots,x_{n})^{\top}\in\mathbb{R}^{n}$ such that for all $i\in N$,
$|a_{i\cdots i}|x_{i}^{m-1}>\sum\limits_{\begin{subarray}{c}i_{2},\ldots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{i_{2}}\cdots x_{i_{m}}.$ (3)
If $\mathcal{A}$ is an $H$-tensor, then $0\notin\sigma(\mathcal{A})$.
$\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$ is $Z$-tensor if all of its off-diagonal entries are non-positive, which is equivalent to write $\mathcal{A}=s\mathcal{I}-\mathcal{B}$, where $s>0$ and $\mathcal{B}$ is a nonnegative tensor. $\mathcal{A}$ is called an $M$-tensor [56] if there exist a nonnegative tensor $\mathcal{B}$ and a positive real number $s>\rho(\mathcal{B})$ such that
$\mathcal{A}=s\mathcal{I}-\mathcal{B}.$
2.2 Preliminaries for tensor quantum information
In quantum information theory [11], a pure quantum state $|v\rangle\in\mathbb{C}^{n}$ is a normal vector, and a mixed state $\rho\in M_{n}(\mathbb{C})$ is a positive semidefinite (PSD) Hermitian matrix with $\operatorname{Tr}(\rho)=1$. A mixed state $\rho\in M_{m}(\mathbb{C})\otimes M_{n}(\mathbb{C})$ is called separable if there exist pure states $\{|w_{i}\rangle\}_{i}\subseteq\mathbb{C}^{m}$ and $\{|v_{i}\rangle\}_{i}\subseteq\mathbb{C}^{n}$ such that
$\rho=\sum\limits_{i}p_{i}|w_{i}\rangle\langle w_{i}|\otimes|v_{i}\rangle\langle v_{i}|,~{}~{}p_{i}\geq 0,~{}~{}\sum p_{i}=1,$
where $\otimes$ denotes the Kronecker product [57], and $\langle v_{i}|$ is the dual (row) vector of $|v_{i}\rangle$. In other words, $\rho$ is separable if and only if it can be written in the form
$\rho=\sum\limits_{i}X_{i}\otimes Y_{i},$
where $X_{i}\in M_{m}(\mathbb{C})$ and $Y_{i}\in M_{n}(\mathbb{C})$ are positive semidefinite matrices. If the state $\rho$ is not separable, it is called an entangled state [58].
Spin-$j$ coherent states $|\alpha\rangle$ is defined for $\alpha=e^{-i\varphi}\cot(\theta/2)$ with $\theta\in[0,\pi]$ and $\varphi\in[0,2\pi[$ by
$|\alpha\rangle=\sum_{l=-j}^{j}\sqrt{\left(\begin{array}[]{c}2j\\ j+l\end{array}\right)}\left[\sin\frac{\theta}{2}\right]^{j-l}\left[\cos\frac{\theta}{2}e^{-i\varphi}\right]^{j+l}|j,l\rangle$ (4)
in the standard angular momentum basis $\{|j,l\rangle:-j\leq$ $l\leq j\}$. The spin coherent state $|\alpha\rangle$ can be considered as a spin-$j$ pointing in the direction $\mathbf{n}=(\sin\theta\cos\varphi,\sin\theta\sin\varphi,\cos\theta)$.
A spin-$j$ state can be seen as a symmetrized state of $2j$ spins $\frac{1}{2}$, and thus it can be written with $m=2j$ [42]. A spin-$j$ coherent state is the tensor product of $2j$ identical spin-$\frac{1}{2}$ coherent states. Since any spin-$\frac{1}{2}$ pure state is a coherent state, spin-$j$ coherent states are by definition the symmetric separable pure states.
A spin-$j$ state $\rho_{\mathcal{C}}$ is classical [38] if and only if it can be decomposed into the sum of spin coherent states with positive weights, that is, there are spin coherent states $|\alpha_{i}\rangle$ such that
$\rho_{\mathcal{C}}=\sum\limits_{i}\omega_{i}|\alpha_{i}\rangle\langle\alpha_{i}|,\quad 0\leq\omega_{i}\leq 1,\sum\limits_{i}\omega_{i}=1.$ (5)
Since a coherent spin state is equivalent to a symmetric separable pure state, the entangled symmetric multiqubit state can not be written as a classical state, that is, it can not be written as in (5). Moreover, The literature on [59] show that classical states are identified with fully separable symmetric states.
3 The tensor-generated matrices and $H$-tensors
In this section, we introduce the concept of tensor-generated matrices. These matrices, derived from tensors, are broadly characterized as the associations between an $m$-th order $n$-dimensional tensor $\mathcal{A}$ and an $n$-dimensional square matrix $A$. This classification streamlines the alignment of elements in tensor $\mathcal{A}$ with their counterparts in matrix $A$, both sharing identical dimensions. We establish that if the tensor-generated matrix $A$ is an $H$-matrix, then the tensor $\mathcal{A}$ must be an $H$-tensor. Moreover, we delve into similar properties shared by the original tensor and its derived matrix, including weak irreducibility and weakly chained diagonal dominance. These methodologies facilitate the extension of nearly all properties of pseudo-diagonal dominance from matrices to higher-order tensors, challenging the prior notion that such properties pertained exclusively to matrices rather than tensors.
For the sake of convenience, the following notation will be used throughout the remainder of this paper:
$s_{ij}(\mathcal{A})=\frac{1}{m-1}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=j,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|,\quad j=1,2,\cdots,n.$
It is important to note that the value of $m$ is greater than or equal to 2. Specifically, when $m=2$, $s_{ij}(A)$ corresponds to the sum of the elements in a row of the matrix minus the diagonal elements.
3.1 The tensor-generated matrices and $H$-tensors
Firstly, we introduce the concept of tensor-generated matrices that characterize as the associations between an $m$-th order $n$-dimensional tensor $\mathcal{A}$ and an $n$-dimensional square matrix $A$.
Definition 1.
Consider $\mathcal{A}\in\mathbb{C}^{[m\times n]}$ as an $m$th-order $n$-dimensional tensor. From this tensor, we derive a square matrix $A=(a_{ij})\in\mathbb{C}^{n,n}$ with $n$ dimensions, defined as:
$a_{ij}=\begin{cases}|a_{ii\cdots i}|-s_{ii}(\mathcal{A}),&i=j,\\ s_{ij}(\mathcal{A}),&i\neq j.\end{cases}$ (6)
This square matrix $A\in\mathbb{C}^{n,n}$ is termed the tensor-generated matrix of tensor $\mathcal{A}$.
In this context, we specifically map the elements of an $m$th-order $n$-dimensional tensor $\mathcal{A}$ onto a square matrix $A$ containing $n$ dimensions. This mapping involves categorizing the $n^{m}$ entries of $\mathcal{A}$ to establish the correspondences between an $m$th-order $n$-dimensional tensor $\mathcal{A}$ and an $n$-dimensional square matrix $A$. Subsequently, we show that if the tensor-generated matrix $A\in\mathbb{C}^{n,n}$ is an $H$-matrix, then the tensor $\mathcal{A}$ is also an $H$-tensor.
Theorem 1.
Let $A\in\mathbb{C}^{n,n}$ be a tensor-generated matrix by tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. If matrix $A$ is an $H$-matrix, then tensor $\mathcal{A}$ is an $H$-tensor.
Proof.
It is a known fact that when the tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$ qualifies as an $H$-tensor, it satisfies the inequality defined in (3),
$|a_{i\cdots i}|x_{i}^{m-1}>\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{i_{2}}\cdots x_{i_{m}}.$
Utilizing the arithmetic-geometric mean inequality, we can represent the right-hand side of the aforementioned inequality as follows:
$\begin{split}&\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{i_{2}}\cdots x_{i_{m}}\\ =&\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|^{\frac{1}{m-1}}x_{i_{2}}\cdots|a_{ii_{2}\cdots i_{m}}|^{\frac{1}{m-1}}x_{i_{m}}\\ \leq&\frac{1}{m-1}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}(|a_{ii_{2}\cdots i_{m}}|x_{i_{2}}^{m-1}+\cdots+|a_{ii_{2}\cdots i_{m}}|x_{i_{m}}^{m-1})\\ =&\frac{1}{m-1}(\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{i_{2}}^{m-1}+\cdots+\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{i_{m}}^{m-1})\\ =&\frac{1}{m-1}\sum\limits_{t=2}^{m}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x^{m-1}_{i_{t}}.\end{split}$ (7)
It is note worthy that the following $m-1$ equalities hold for $p=2,\cdots,m$
$\begin{split}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}&|a_{ii_{2}\cdots i_{m}}|x_{i_{p}}^{m-1}=\sum\limits_{\begin{subarray}{c}i_{k}\in N,i_{p}=1,k=2\cdots m,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{1}^{m-1}\\ &+\cdots+\sum\limits_{\begin{subarray}{c}i_{k}\in N,i_{p}=n,k=2\cdots m,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{n}^{m-1},\end{split}$ (8)
If we set $p=2$, then for $m=2$, we obtain a matrix where the indices $i_{3}\ldots i_{m}$ are eliminated. In this scenario, there is
$\sum\limits_{\delta_{ii_{2}}=0}|a_{ii_{2}}|x_{i_{2}}=\sum\limits_{\delta_{ii_{2}}=0}|a_{i1}|x_{1}+\cdots+\sum\limits_{\delta_{ii_{2}}=0}|a_{in}|x_{n}.$
We sum up the $q$-th term from all the aforementioned equalities in (8) for $q=1$ to $n$,
$\begin{split}\sum\limits_{\begin{subarray}{c}i_{k}\in N,i_{2}=q,k=2\cdots m,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}&|a_{ii_{2}\cdots i_{m}}|x_{q}^{m-1}+\cdots+\sum\limits_{\begin{subarray}{c}i_{k}\in N,i_{m}=q,k=2\cdots m,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{q}^{m-1}\\ =&\sum\limits_{k=2}^{n}\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=q,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{q}^{m-1}.\end{split}$
It follows that
$\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}\sum\limits_{t=2}^{m}|a_{ii_{2}\cdots i_{m}}|x^{m-1}_{i_{t}}=\sum\limits_{q=1}^{n}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=q,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{q}^{m-1}.$ (9)
Based on equations (7) and (9), it can be inferred that
$\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{i_{2}}\cdots x_{i_{m}}\leq\frac{1}{m-1}\sum\limits_{t=2}^{m}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x^{m-1}_{i_{t}}=\sum\limits_{j=1}^{n}s_{ij}(\mathcal{A})x_{j}^{m-1}.$ (10)
In addition, we construct the matrix $B=(b_{ij})=AD\in\mathbb{C}^{n,n}$ as follows;
$b_{ij}=\begin{cases}(|a_{ii\cdots i}|-s_{ii}(\mathcal{A}))x_{i}^{m-1},&i=j,\\ s_{ij}(\mathcal{A})x_{j}^{m-1},&i\neq j,\end{cases}$ (11)
where $D=diag(x^{m-1}_{1},x^{m-1}_{2},\cdots,x^{m-1}_{n})$.
Given that the tensor-generated matrix $A$ is an $H$-matrix, there exists a positive diagonal matrix $D=\text{diag}(x^{m-1}_{1},x^{m-1}_{2},\ldots,x^{m-1}_{n})$ such that $B=AD$ forms a strictly diagonally dominant matrix. Hence, the following inequality holds:
$(|a_{ii\cdots i}|-s_{ii}(\mathcal{A}))x_{i}^{m-1}>\sum\limits_{j\neq i}^{n}s_{ij}(\mathcal{A})x_{j}^{m-1}.$
Subsequently, the following inequalities are derived using (10):
$\begin{split}|a_{ii\cdots i}|x_{i}^{m-1}>&\sum\limits_{j=1}^{n}s_{ij}(\mathcal{A})x_{j}^{m-1}\\ =&\frac{1}{m-1}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}\sum\limits_{t=2}^{m}|a_{ii_{2}\cdots i_{m}}|x^{m-1}_{i_{t}}\\ \geq&\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|x_{i_{2}}\cdots x_{i_{m}}.\end{split}$
Consequently, tensor $\mathcal{A}$ qualifies as an $H$-tensor. Thus, the proof is concluded. ∎
Remark 1.
It is important to note that when $x_{i_{2}}=\cdots=x_{i_{m}}=1$, i.e., when the equality in (7) is satisfied, there exist
$\begin{split}r_{i}(\mathcal{A})=&\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|=\frac{1}{m-1}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}\sum\limits_{t=2}^{m}|a_{ii_{2}\cdots i_{m}}|\\ =&\frac{1}{m-1}\sum\limits_{j=1}^{n}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=j,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|\\ =&s_{ii}(\mathcal{A})+\sum\limits_{j=1,j\neq i}^{n}s_{ij}(\mathcal{A}).\end{split}$ (12)
This demonstrates that Definition 1 extends the concept of $r_{i}(\mathcal{A})$, representing a particular case within our framework.
Definition 2.
[56, Definition 3.14] A tensor $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$ is called a (strictly) diagonally dominant tensor if for $i\in N$,
$|a_{i\cdots i}|\geq(>)r_{i}(\mathcal{A})=\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|.$ (13)
And we denote
[…]
Given the construction of the tensor-generated matrix $A\in\mathbb{C}^{n,n}$ from the tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$, the following conclusion can be inferred.
Theorem 2.
Let $A\in\mathbb{C}^{n,n}$ be a tensor-generated matrix by tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. If the tensor-generated matrix $A$ is (strictly) diagonally dominant matrix, then tensor $\mathcal{A}$ is (strictly) diagonally dominant tensor.
Proof.
If a tensor-generated matrix $A\in\mathbb{C}^{n,n}$ is derived from tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$ and exhibits (strict) diagonal dominance, then the inequality
$|a_{ii\cdots i}|-s_{ii}(\mathcal{A})\geq(>)\sum\limits_{j=1,j\neq i}^{n}s_{ij}(\mathcal{A})$
leads to the conclusion derived from (12):
$|a_{ii\cdots i}|\geq(>)\sum\limits_{j=1}^{n}s_{ij}(\mathcal{A})=r_{i}(\mathcal{A}).$
Thus, the proof is concluded. ∎
Remark 2.
If there exists some $i_{0}\in N$ such that $|a_{i_{0}i_{0}\cdots i_{0}}|-s_{i_{0}i_{0}}(\mathcal{A})\leq 0$, it is evident that $\mathcal{A}$ is not a diagonally dominant tensor. Therefore, without special explanation, we always discuss diagonal dominance of tensors in the case of $|a_{ii\cdots i}|>s_{ii}(\mathcal{A})$ throughout this paper.
Remark 3.
As stated in equation (12), a tensor satisfying diagonal dominance, as defined in Definition 2, holds true when $|a_{ii\cdots i}|>s_{ii}(\mathcal{A})$.
According to Theorems 1 and 2, the study of the class of $H$-tensors can be approached by considering the tensor-generated matrix as a diagonally dominant matrix or an $H$-matrix.
Given an $n$-dimensional the tensor-generated matrix $A=(a_{ij})\in\mathbb{C}^{n\times n}$, let us define:
$P_{i}(A)=\sum\limits_{j=1,j\neq i}^{n}s_{ij}(\mathcal{A})=\frac{1}{m-1}\sum\limits_{j\neq i}^{n}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=j,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|,$
$Q_{i}(A)=\sum\limits_{i=1,i\neq j}^{n}s_{ij}(\mathcal{A})=\frac{1}{m-1}\sum\limits_{i\neq j}^{n}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=j,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|.$
Various forms of diagonal dominance exist for matrices. Here, we exemplify the fundamental types of diagonally dominant matrices.
For a matrix $A=(a_{ij})\in\mathbb{C}^{n\times n}$, it is defined as diagonally dominant ($D_{n}$) if, for each $i\in N$,
$|a_{ii}|\geq P_{i}(A).$ (14)
Matrix $A=(a_{ij})\in\mathbb{C}^{n\times n}$ is termed a doubly diagonally dominant matrix ($DD_{n}$) if, for all $i,j\in N$, $i\neq j$,
$|a_{ii}||a_{jj}|\geq\big{(}P_{i}(A)\big{)}\big{(}P_{j}(A)\big{)}.$ (15)
A matrix $A=(a_{ij})\in\mathbb{C}^{n\times n}$ is considered a $\gamma$-diagonally dominant matrix ($D^{\gamma}_{n}$) if there exists $\gamma\in[0,1]$ such that
$|a_{ii}|\geq\gamma\big{(}P_{i}(A)\big{)}+(1-\gamma)\big{(}Q_{i}(A)\big{)},\quad\forall i\in N.$ (16)
Matrix $A=(a_{ij})\in\mathbb{C}^{n\times n}$ is termed a product $\gamma$-diagonally dominant matrix ($PD^{\gamma}_{n}$) if there exists $\gamma\in[0,1]$ such that
$|a_{ii}|\geq\big{(}P_{i}(A)\big{)}^{\gamma}\big{(}Q_{i}(A)\big{)}^{1-\gamma},\quad\forall i\in N.$ (17)
When all inequalities in (14)-(17) hold strictly, $A$ is categorized as a strictly diagonally dominant matrix ($SD_{n}$), strictly doubly diagonally dominant matrix ($SDD_{n}$), strictly $\gamma$-diagonally dominant matrix ($SD^{\gamma}_{n}$), and strictly product $\gamma$-diagonally dominant matrix ($SPD^{\gamma}_{n}$), respectively. Furthermore, $A$ is referred to as a generalized diagonally dominant matrix ($GSD_{n}$), generalized $\gamma$-diagonally dominant matrix ($GSD^{\gamma}_{n}$), and generalized product $\gamma$-diagonally dominant matrix ($GSPD^{\gamma}_{n}$) if there exists a positive diagonal matrix $D$ such that $AD$ is a strictly diagonally dominant matrix, strictly $\gamma$-diagonally dominant matrix, and strictly product $\gamma$-diagonally dominant matrix, respectively.
It is well known that a matrix $A$ is a nonsingular $H$-matrix if there exists a positive diagonal matrix $D=(x_{1},\cdots,x_{n})^{\top}$ ($x_{i}>0$) for $i\in N$ such that $AD$ is a strictly diagonally dominant matrix ($SD_{n}$) [60].
Lemma 1.
[57, 60] Let $A\in\mathbb{C}^{n,n}$. If matrix $A$ meets one of the following conditions, then matrix $A$ is a $H$-matrix.
(1) $A$ is a strictly diagonally dominant matrix ($SD_{n}$).
(2) $A$ is strictly doubly diagonally dominant matrix ($SDD_{n}$).
(3) $A$ is a strictly $\gamma$-diagonally dominant matrix ($SD^{\gamma}_{n}$).
(4) $A$ is a strictly product $\gamma$-diagonally dominant matrix ($SPD^{\gamma}_{n}$).
(5) $A$ is a generalized diagonally dominant matrix ($GSD_{n}$), generalized $\gamma$-diagonally dominant matrix ($GSD^{\gamma}_{n}$), and generalized product $\gamma$-diagonally dominant matrix ($GSPD^{\gamma}_{n}$).
Nonetheless, properties beyond diagonal dominance do not straightforwardly extend to higher-order tensors due to their exclusive applicability to matrices.
However, a connection can be established between these diagonally dominant properties and higher-order tensors through the tensor-generated matrix, facilitating the alignment of $H$-tensor and $H$-matrix evaluations, thereby enhancing $H$-tensor theory.
Likewise, we can illustrate that the subsequent classical matrix outcome extends to higher-order tensors through the tensor-generated matrix in conjunction with Lemma 1.
Theorem 3.
Consider a tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$ that generates a matrix $A\in\mathbb{C}^{n,n}$. If the tensor-generated matrix $A$ satisfies any of the conditions outlined in Lemma 1, it implies that the tensor $\mathcal{A}$ is an $H$-tensor.
It is crucial to highlight that extending fundamental diagonally dominant properties to higher-order tensors via the tensor-generated matrix is feasible. Theorem 3 illustrates the extension of the remaining properties in Lemma 1 to higher-order tensors, resembling property (1) in Lemma 1. This establishes the theoretical foundation for the content in the fifth section.
3.2 The tensor-generated matrix and weakly irreducible diagonally dominant tensor
Emphasizing that it is feasible to extend fundamental diagonally dominant properties to higher-order tensors via the tensor-generated matrix is crucial. Most of the determination findings for $H$-matrices can be extended to higher-order tensors, including weakly irreducible diagonally dominant tensors and weakly chained diagonally dominant tensors.
Definition 3.
[61, Definition 2.1] Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$. Then $\mathcal{A}$ is called reducible, if there exists a nonempty proper index $I\subset N$ such that
$a_{i_{1}i_{2}\cdots i_{m}}=0,\quad\forall i_{1}\in I,\quad\forall i_{2},\cdots,i_{m}\notin I.$
If $\mathcal{A}$ is not reducible, then $\mathcal{A}$ is called irreducible.
Definition 4.
[62, Definition 2.2] Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$ and $R(|\mathcal{A}|)$, the representation of $\mathcal{A}$, be a $n\times n$ matrix whose $(i,j)$-th entry is given by
$R(|\mathcal{A}|)_{ij}=\sum\limits_{j\in\{i_{2}\cdots i_{m}\}}|a_{ii_{2}\cdots i_{m}}|.$
Tensor $\mathcal{A}$ is called weakly reducible if $R(|\mathcal{A}|)$ is a reducible matrix. Otherwise, tensor $\mathcal{A}$ is weakly irreducible
There is another equivalent about weakly irreducible tensor. An order $m$ dimension $n$ tensor $\mathcal{A}$ is called weakly reducible, if there exists a nonempty proper index subset $I\subset N$ such that
$a_{i_{1}i_{2}\cdots i_{m}}=0,\quad\forall i_{1}\in I,\quad\exists i_{j}\notin I,\quad j=2,\cdots,m.$
If $\mathcal{A}$ is not weakly reducible, then we call $\mathcal{A}$ weakly irreducible [62, 63]. In the context of matrices, it is worth noting that there is no distinction between irreducible and weakly irreducible matrices. However, while an irreducible tensor must also be weakly irreducible, the reverse is not necessarily true.
Lemma 2.
[63, Theorem 3.4] Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$. If $\mathcal{A}$ is irreducible,
$|a_{i\cdots i}|\geq r_{i}(\mathcal{A}),\quad\forall i\in N,$
and strictly inequality holds for at least one $i$, then $\mathcal{A}$ is an $H$-tensor.
It can be demonstrated that the tensor-generated matrix $A$ and tensor $\mathcal{A}$ exhibit equivalent weak irreducibility, facilitating the extension of conclusions concerning $H$-matrices from irreducibility to higher-order tensors.
Theorem 4.
Consider $A\in\mathbb{C}^{n,n}$ as the matrix generated by the tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. Then the tensor $\mathcal{A}$ is weakly irreducible if and only if the tensor-generated matrix $A$ is irreducible.
Proof.
Necessity: Assume that $\mathcal{A}$ is a weakly irreducible tensor. Suppose the tensor-generated matrix $A$ is reducible. Then there exists a nonempty proper index subset $I\subset N$ such that $a_{ij}=0,\forall i\in I,\forall j\notin I$, meaning that we have
$a_{ij}=s_{ij}(\mathcal{A})=\frac{1}{m-1}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{k-1},i_{k+1},\cdots,i_{m}\in N,i_{k}=j,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|=0,j\neq i.$
Additionally, we find $a_{ii_{2}\cdots i_{m}}=0,\forall i\in I,\exists i_{2}\cdots i_{m}\notin I$, implying that $\mathcal{A}$ is a weakly reducible tensor, leading to a contradiction.
Sufficiency: Given that the tensor-generated matrix $A$ is irreducible. Assuming that the tensor $\mathcal{A}$ is weakly reducible, there exists a nonempty proper index subset $I\subset N$ such that $a_{ii_{2}\cdots i_{m}}=0,\forall i\in I,\exists i_{2}\cdots i_{m}\notin I$, which implies
$s_{ij}(\mathcal{A})=\frac{1}{m-1}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{k-1},i_{k+1},\cdots,i_{m}\in N,i_{k}=j,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|=0,j\neq i.$
Equivalently, there exists a nonempty proper index subset $I\subset N$ such that $a_{ij}=s_{ij}(\mathcal{A})=0,\forall i\in I,\forall j\notin I$, contradicting the irreducibility of matrix $A$. This concludes the proof. ∎
Corollary 1.
Consider $A\in\mathbb{C}^{n,n}$ as a matrix generated by the tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. If the tensor $\mathcal{A}$ is irreducible, then the tensor-generated matrix $A$ is also irreducible.
Leveraging Theorem 2, Theorem 4, and Lemma 2, we can draw the subsequent deductions.
Theorem 5.
Let $A\in\mathbb{C}^{n,n}$ be a tensor-generated matrix by tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. If the tensor-generated matrix $A$ is irreducible diagonally dominant matrix and strictly inequality holds for at least one, then $\mathcal{A}$ is weakly irreducible diagonally dominant tensor. Moreover, tensor $\mathcal{A}$ is $H$-tensor.
Proof.
If the tensor-generated matrix $A\in\mathbb{C}^{n,n}$ by tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$ is irreducible, then Theorem 4 implies that tensor $A$ is weakly irreducible. Since the tensor-generated matrix $A\in\mathbb{C}^{n,n}$ is diagonally dominant, according to Theorem 2, tensor $\mathcal{A}$ is diagonally dominant. Furthermore, based on Lemma 2, tensor $\mathcal{A}$ is an $H$-tensor. The proof is completed. ∎
Remark 4.
Since tensor $\mathcal{A}$ is weakly irreducible if and only if tensor-generated matrix $A\in\mathbb{C}^{n,n}$ by tensor $\mathcal{A}$ is irreducible. This aligns with the finding from (Theorem 3.4 in [63]) that a weakly irreducible diagonally dominant tensor is an $H$-tensor.
Corollary 2.
Consider $A\in\mathbb{C}^{n,n}$ as a tensor-generated matrix by irreducible tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. If the tensor-generated matrix $A$ is diagonally dominant matrix and strictly inequality holds for at least one, then $\mathcal{A}$ is $H$-tensor.
Theorem 4 demonstrates that both the tensor and its tensor-generated matrix demonstrate equivalent weak irreducibility. This observation directs our focus towards exploring the properties of $H$-tensors in the realm of weak irreducibility.
Lemma 3.
[57, 60] If matrix $A\in\mathbb{C}^{n,n}$ meets one of the following conditions, then matrix $A$ is a $H$-matrix.
(1) $A$ is a irreducible diagonally dominant matrix and strictly inequality holds for at least one.
(2) $A$ is a irreducible doubly diagonally dominant matrix and strictly inequality holds for at least one.
(3) $A$ is a irreducible $\gamma$-diagonally dominant matrix and strictly inequality holds for at least one.
(4) $A$ is a irreducible product $\gamma$-diagonally dominant matrix and strictly inequality holds for at least one.
Theorem 5 affirms the feasibility of extending from case (1) in Lemma 3 to higher-order tensors. This extension is not limited to that particular case alone; indeed, all scenarios outlined in Lemma 3 are generalizable to the tensor scenario through the utilization of the tensor-generated matrix.
In essence, the deductions stemming from Lemma 3 and Theorem 1 lead us to the following conclusions. The proof methodology mirrors a comparable approach, and we will avoid redundancy by omitting its repetition.
Theorem 6.
Consider $A\in\mathbb{C}^{n,n}$ as a tensor-generated matrix by tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. If matrix $A$ satisfies one of the conditions in Lemma 3, then $\mathcal{A}$ is an $H$-tensor.
3.3 The tensor-generated matrix and weakly chained diagonally dominant tensor
It is crucial to emphasize the possibility of extending fundamental diagonally dominant properties to higher-order tensors through the utilization of the tensor-generated matrix. Most determinant findings for $H$-matrices are transferable to higher-order tensors, encompassing weakly chained diagonally dominant tensors. Prior to outlining our conclusions, it is necessary to introduce pertinent findings within the directed graph linked to tensors.
Definition 5.
[…]
Lemma 4.
[…]
Definition 6.
[64, Definition 15] A tensor $\mathcal{A}$ is weakly chained diagonally dominant if all of the following are satisfied: (i) $\mathcal{A}$ is diagonally dominant. (ii) $J(\mathcal{A})$ is nonempty. (iii) For each $i^{1}\notin J(\mathcal{A})$, there exists a walk $i^{1}\rightarrow i^{2}\rightarrow\cdots\rightarrow i^{k}$ in graph $\mathcal{A}$ such that $i^{k}\in J(\mathcal{A})$.
Lemma 5.
[64, Theorem 1] Let $\mathcal{A}$ be a weakly chained diagonally dominant $Z$-tensor with nonnegative diagonals, then $\mathcal{A}$ is a $M$-tensor.
It can be demonstrated that when tensor $\mathcal{A}$ exhibits weakly chained diagonally dominant characteristics, the corresponding tensor-generated matrix $A$ is diagonally dominant. Moreover, a non-zero element chain exists from $i\notin J(\mathcal{A})$ to $j\in J(\mathcal{A})$ for $i,j\in N$ (where $i\neq j$).
Theorem 7.
Let $A\in\mathbb{C}^{n,n}$ be a tensor-generated matrix by tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. If the tensor-generated matrix $A$ is a diagonally dominant matrix with a non-zero element chain from $i\notin J(\mathcal{A})$ to $j\in J(\mathcal{A})$ for $i,j\in N$, then tensor $\mathcal{A}$ is a weakly chained diagonally dominant tensor. Furthermore, tensor $\mathcal{A}$ is an $H$-tensor.
Proof.
Drawing upon Theorem 2, we promptly deduce (i) and (ii) from Definition 6. Furthermore, according to Lemma 4, if the tensor-generated matrix $A$ has a non-zero element chain from $i\notin J(A)$ to $j\in J(A)$ for $i,j\in N$, then $\mathcal{A}$ also features a path $i^{1}\rightarrow i^{2}\rightarrow\cdots\rightarrow i^{k}$ within the graph $\mathcal{A}$ such that $i^{k}\in J(\mathcal{A})$, where $i^{1}=i$ and $i^{k}=j$. Consequently, tensor $\mathcal{A}$ aligns with the definition of weakly chained diagonally dominant tensors as per Definition 6. By virtue of Lemma 5, tensor $\mathcal{A}$ qualifies as an $H$-tensor. This concludes the proof. ∎
In summary, when faced with strict inequality, weak irreducibility, or weak chains, we have the ability to convert the problem of determining an $H$-tensor into the problem of determining an $H$-matrix. This conversion provides us with the advantage of leveraging the well-established $H$-matrix theory to address more intricate problems involving $H$-tensors. As a result, the development of $H$-tensor theory is significantly advanced.
4 Application in Classicality for spin states
This section delves into the application of tensor-generated matrices in classical spin states. Leveraging the spectral properties of $H$-tensors from Section 3, we can establish adequate criteria for positive definite tensors. Utilizing the tensor representation of spin states and tensor-generated matrices in symmetric scenarios, we introduce criteria for separability (classicality) in spin-$j$ states.
4.1 Classicality for spin states with symmetric coefficient tensors
The literature referenced in [26] introduces a tensorial representation for spin states, expanding a spin-$j$ density matrix $\rho$ as follows: The $4^{m}$ Hermitian matrices $S_{\mu_{1}\ldots\mu_{m}}$ (with $m=2j$ ) is defined as
$S_{\mu_{1}\cdots\mu_{m}}=P\left(\sigma_{\mu_{1}}\otimes\sigma_{\mu_{2}}\cdots\otimes\sigma_{\mu_{m}}\right)P^{\dagger},\quad 0\leqslant\mu_{i}\leqslant 3,$
with $P$ being the projector onto the symmetric subspace of tensor products of $N$ spins- $\frac{1}{2}$ (the subspace spanned by Dicke states). Here, $\sigma_{a},1\leqslant a\leqslant 3$, be the usual Pauli matrices and $\sigma_{0}$ be the 2×2$\times$2 × 2 identity matrix as follows:
$\sigma_{1}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\sigma_{2}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix},\sigma_{3}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},\sigma_{0}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}.$
For a general spin-$j$, the $S_{\mu_{1}\mu_{2}\cdots\mu_{m}}$ provide an overcomplete basis over which $\rho$ can be expanded, that is, any state can be expressed as
$\rho=\frac{1}{2^{m}}\mathcal{A}_{\mu_{1}\mu_{2}\cdots\mu_{N}}S_{\mu_{1}\mu_{2}\cdots\mu_{m}},$ (18)
with coefficients
$\mathcal{A}_{\mu_{1}\mu_{2}\cdots\mu_{m}}=tr(\rho S_{\mu_{1}\mu_{2}\cdots\mu_{m}}),$ (19)
real and invariant under permutation of the indices. These $\mathcal{A}_{\mu_{1}\mu_{2}\cdots\mu_{m}}$ can be viewed as a symmetric order-$N$ tensor. Thus, we denote Equation (19) as the tensor representation of $\rho$.
The tensor corresponding to a spin coherent state $|\alpha\rangle$ aligned with the direction n is represented as
$\mathcal{A}_{\mu_{1}\mu_{2}\cdots\mu_{m}}=\langle\alpha|S_{\mu_{1}\mu_{2}\cdots\mu_{m}}|\alpha\rangle=n_{\mu_{1}}n_{\mu_{2}}\cdots n_{\mu_{m}},$
where $n_{0}=1$ and $\textbf{n}=(n_{1},n_{2},n_{3})$.
It is worth noting that the generalized Bloch representation (18) shares several crucial properties with the Bloch representation of a spin-$\frac{1}{2}$. From (19), we observe that the coordinates of a coherent state are derived simply by multiplying the components of the 4-vector $n_{\mu}^{(i)}=(1,\textbf{n}^{(i)})$. This extends the concept that the Bloch vector representing a spin-$\frac{1}{2}$ state aligns with the directions defined by the angles of the coherent state. Consequently, we have $\mathbf{n}=(\sin\theta\cos\varphi,\sin\theta\sin\varphi,\cos\theta)\geq 0$, where $\theta\in[0,\frac{\pi}{2}]$ and $\varphi\in[0,\frac{\pi}{2}]$.
The concept of classicality defined in (5) can be reformulated using tensors. A state is deemed classical if and only if there exist positive weights $\omega_{i}$ and unit vectors $\textbf{n}^{(i)}$ such that its tensor of coordinates $\mathcal{A}$ can be expressed as
$\mathcal{A}=\sum\limits_{i}\omega_{i}n_{\mu_{1}}^{(i)}n_{\mu_{2}}^{(i)}\cdots n_{\mu_{m}}^{(i)}=\sum\limits_{i=1}^{r}\omega_{i}\big{(}n_{\mu}^{(i)}\big{)}^{\otimes m},$ (20)
where $n_{\mu}^{(i)}=(1,\textbf{n}^{(i)})$. It is evident that $\mathcal{A}$ is a symmetric tensor of rank $r$.
Within Section 3, when strict inequality, weak irreducibility, or weak chain conditions are present, the calculation of the $H$-tensor can be streamlined to computing the $H$-matrix. This streamlined approach allows us to apply the well-established $H$-matrix theory to address more intricate $H$-tensor challenges, thereby significantly advancing the realm of $H$-tensor theory. Notably, we can leverage the spectral characteristics of $H$-tensors to establish sufficient conditions for positive semidefinite tensors.
Lemma 6.
Consider $\mathcal{A}\in\mathbb{C}^{[m\times n]}$ as an even-order real symmetric tensor where $a_{kk\cdots k}\geq 0$ for all $k\in\mathbb{N}$, and let $A\in\mathbb{C}^{n,n}$ represent the tensor-generated matrix by the tensor $\mathcal{A}$. If matrix $A$ qualifies as an $H$-matrix, it implies that the tensor $\mathcal{A}$ is positive semidefinite.
In this context, we introduce the criterion for determining the classicality of spin states, drawing from Lemma 6 and the relationship established in Section 3.
Theorem 8.
For a spin-$j$ state, if its representation tensor $\mathcal{A}_{\mu_{1}\cdots\mu_{2j}}\in\mathbb{R}^{[2j\times 4]}$ is a real tensor with $a_{kk\cdots k}\geq 0$ for all $k\in\{1,2,3,4\}$, and $A\in\mathbb{C}^{4,4}$ is a tensor-generated matrix by tensor $\mathcal{A}$. If matrix $A$ is a $H$-matrix, the tensor $\mathcal{A}$ is positive semidefinite. In other words, spin-$j$ state is classical.
Proof.
Given a spin-$j$ state $\rho$, there exists a corresponding coefficient tensor $\mathcal{A}_{\mu_{1}\mu_{2}\cdots\mu_{m}}=tr(\rho S_{\mu_{1}\mu_{2}\cdots\mu_{m}})$ as shown in (19). A spin-$j$ coherent state can be seen as a symmetrized tensor product of $2j$ spins $\frac{1}{2}$ coherent states, and thus it can be written with representation tensor $\mathcal{A}_{\mu_{1}\cdots\mu_{2j}}\in\mathbb{R}^{[2j\times 4]}.$
If the representation tensor $\mathcal{A}_{\mu_{1}\cdots\mu_{2j}}\in\mathbb{R}^{[2j\times 4]}$ is a real tensor with $a_{kk\cdots k}\geq 0$ for all $k\in\{1,2,3,4\}$, and the resulting tensor-generated matrix $A\in\mathbb{C}^{4,4}$ is an $H$-matrix, then by Lemma 6, we can conclude that the representation tensor $\mathcal{A}_{\mu_{1}\cdots\mu_{2j}}\in\mathbb{R}^{[2j\times 4]}$ is positive semidefinite.
When $j$ is an integer (thus $m=2j$ is even), the right-hand side of equation (20) is consistently positive. For any $x\in\mathbb{R}^{n}$, it follows that
$\mathcal{A}x^{m}=\sum\limits_{i}\omega_{i}n_{\mu_{1}}^{(i)}n_{\mu_{2}}^{(i)}\cdots n_{\mu_{m}}^{(i)}x_{\mu_{1}}x_{\mu_{2}}\cdots x_{\mu_{m}}=\sum\limits_{i=1}^{r}\omega_{i}\left[\left(n_{\mu}^{(i)}\right)^{\top}x\right]^{\otimes m}\geq 0,$
illustrating the positive semidefiniteness of the even-order tensor $\mathcal{A}$. Consequently, we have demonstrated that the spin-$j$ state adheres to classicality under condition (20). The proof is completed. ∎
It is worth noting that when $j$ is an integer, all expansion coefficients correspond to tensors that are fourth-order tensors of even order. This not only establishes a theoretical foundation for deriving the aforementioned criterion using even-order semi-positive definite tensors but also preserves the physical significance of all tensor elements. Specifically, the components of all 4-dimensional vectors $n_{\mu}^{(i)}=(1,\textbf{n}^{(i)})$ equate to 1 and correspond to the polar coordinates of points $\textbf{n}^{(i)}$ on the Bloch sphere.
4.2 Classicality for spin states with strongly symmetric coefficient tensors
Quantum tasks are significantly streamlined in symmetric scenarios due to the restriction of optimization to symmetrically separable states. Yet, dealing with the complexity of multipartite states remains a formidable challenge. Addressing these challenges requires exploring quantum states existing within a smaller space compared to symmetric states. A recent contribution by Xiong et al. [65] introduces the notion of multipartite strongly symmetric states, representing a subspace within symmetric states.
Definition 7.
[65, 66] For $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{R}^{[m\times n]}$, we denote
$\mathcal{N}=\{(i_{1}i_{2}\cdots i_{m}):1\leq i_{k}\leq n,k=1,2,\cdots,m\}.$
For $(i_{1}i_{2}\cdots i_{m})\in\mathcal{N}$, denote by $[(i_{1}i_{2}\cdots i_{m})]$ the set of all distinct members in $\{i_{1}i_{2}\cdots i_{m}\}$. Let $(i_{1}i_{2}\cdots i_{m})\in\mathcal{N}$, $(j_{1}j_{2}\cdots j_{m})\in\mathcal{N}$. We say that $(i_{1}i_{2}\cdots i_{m})$ is similar to $(j_{1}j_{2}\cdots j_{m})$, if $[(i_{1}i_{2}\cdots i_{m})]=[(j_{1}j_{2}\cdots j_{m})]$, denote by $(i_{1}i_{2}\cdots i_{m})\sim(j_{1}j_{2}\cdots j_{m})$. We call $\mathcal{A}$ is strongly symmetric, if for any $(i_{1}i_{2}\cdots i_{m}),(j_{1}j_{2}\cdots j_{m})\in\mathcal{N}$, $(i_{1}i_{2}\cdots i_{m})$ $\sim(j_{1}j_{2}\cdots j_{m})$, it holds that $a_{i_{1}i_{2}\cdots i_{m}}=a_{j_{1}j_{2}\cdots j_{m}}$. If $m$-multipartite state $|\Phi\rangle$ is corresponding to a strongly symmetric tensor $\mathcal{A}_{|\Phi\rangle}$. We denote all $m$-multipartite strongly symmetric state of $n$ qudits as $SS^{m}(\mathbb{R}^{n})$.
It is important to highlight that a real strongly symmetric tensor is symmetric, with the exception of when $m=2$. Subsequently, we demonstrate that the symmetry observed in the tensor-generated matrix from the tensor aligns with that strong symmetry of the original tensor.
Theorem 9.
Let $A\in\mathbb{C}^{n,n}$ be a tensor-generated matrix by tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$. If the tensor $\mathcal{A}$ is strongly symmetric, if and only if the tensor-generated matrix $A$ is symmetric.
Proof.
Necessity: Since $\mathcal{A}$ is a symmetric tensor. Then it holds that $a_{i_{1}i_{2}\cdots i_{m}}=a_{j_{1}j_{2}\cdots j_{m}}$ for $(i_{1}i_{2}\cdots i_{m})$, $(j_{1}j_{2}\cdots j_{m})\in\mathcal{N}$, $(i_{1}i_{2}\cdots i_{m})=\pi(j_{1}j_{2}\cdots j_{m})$. For the tensor-generated matrix $A\in\mathbb{C}^{n,n}$, it follows that
$\begin{split}s_{ij}(\mathcal{A})&=\frac{1}{m-1}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=j,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|\\ &=\frac{1}{m-1}\sum\limits_{k=2}^{m}\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=i,\\ \delta_{ji_{2}\cdots i_{m}}=0\end{subarray}}|a_{ji_{2}\cdots i_{m}}|\\ &=s_{ji}(\mathcal{A}).\end{split}$
This shows that the tensor-generated matrix $A$ is symmetric.
Sufficiency: The tensor-generated matrix $A\in\mathbb{C}^{n,n}$ is symmetric, yields
$\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=j,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|=\sum\limits_{\begin{subarray}{c}i_{t}\in N,t\in N\setminus\{1,k\},i_{k}=i,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ji_{2}\cdots i_{m}}|.$
In other word, one has
$|a_{ii_{2}\cdots i_{k-1}ji_{k+1}\cdots i_{m}}|=|a_{ji_{2}\cdots i_{k-1}ii_{k+1}\cdots i_{m}}|,i_{t}\in N,t\in N\setminus\{1,k\},$
which indicates that tensor $\mathcal{A}$ is strong symmetric. The proof is completed. ∎
Subsequently, we introduce the criterion for determining the classicality of spin states with strongly symmetric coefficient tensors.
Theorem 10.
For a spin-$j$ state, if its representation tensor $\mathcal{A}_{\mu_{1}\cdots\mu_{2j}}\in\mathbb{R}^{[2j\times 4]}$ is a strongly symmetric tensor, and tensor-generated matrix $A\in\mathbb{C}^{4,4}$ is a $H$-tensor with $|a_{ii\cdots i}|\geq s_{ii}(\mathcal{A})$. In this situation, spin-$j$ state is symmetric classical.
Proof.
According to Theorem 9, the tensor-generated matrix $A\in\mathbb{C}^{4,4}$ exhibits symmetry. If the tensor-generated matrix $A\in\mathbb{C}^{4,4}$ is an $H$-tensor with $|a_{ii\cdots i}|\geq s_{ii}(\mathcal{A})$, then the matrix $A$ possesses nonnegative diagonal elements, indicating its positive semidefiniteness. Drawing on Lemma 6, we deduce that the representation tensor $\mathcal{A}_{\mu_{1}\cdots\mu_{2j}}\in\mathbb{R}^{[2j\times 4]}$ is positive semidefinite. Consequently, the spin-$j$ state demonstrates symmetric classical behavior. ∎
The authors in [66] demonstrate that a strongly symmetric hierarchically dominated nonnegative tensor is equivalent to a completely positive tensor. Moreover, they specify that a completely positive tensor of even order $m$ is a positive semidefinite tensor, i.e.,
$\mathcal{A}x^{m}=\sum_{k=1}^{r}\left(u^{(k)}\right)^{\otimes m}x^{m}=\sum_{k=1}^{r}\left[\left(u^{(k)}\right)^{\top}x\right]^{\otimes m}\geq 0,\quad\forall x\in\mathbb{R}^{n}.$
In essence, determining that the coefficient tensor $\mathcal{A}$ of a strongly symmetric order-$m$ is either a strongly symmetric hierarchically dominated nonnegative tensor or a completely positive tensor implies the classical nature of the spin-$j$ state.
Corollary 3.
For a spin-$j$ state, if its representation tensor $\mathcal{A}_{\mu_{1}\cdots\mu_{2j}}\in\mathbb{R}^{[2j\times 4]}$ is a strongly symmetric hierarchically dominated nonnegative tensor, then tensor-generated matrix $A\in\mathbb{C}^{4,4}$ is a symmetric $H$-tensor and a copositive matrix. This shows spin-$j$ state is symmetric classical, i.e,
$\mathcal{A}=\sum_{k=1}^{r}\left(u^{(k)}\right)^{\otimes m}.$
In this section, we employed the matrix generated by tensors to analyze the classical characteristics of spin-$j$ states. We investigated the characteristics of the coefficient tensor under both symmetric and strongly symmetric conditions. Furthermore, we elucidated a direct correlation between the strong symmetry of the tensor and the symmetry of the tensor-generating matrix.
5 Application in $H$-Eigenvalue inclusion sets for high order tensors
In this section, the focus is on the inclusion sets of $H$-Eigenvalues for higher-order tensors. Traditional eigenvalue localization results for matrices, including Brauer’s Ovals of Cassini sets, Ostrowski sets, and $S$-type sets, do not directly apply to tensor eigenvalues. To address this challenge, an effort is made to establish a connection between eigenvalue localization for higher-order tensors and tensor-generated matrices, expanding on the connection established in Section 3.
By leveraging this connection, the distribution of $H$-eigenvalues for higher-order tensors is determined by linking it to the distribution of eigenvalues for the subset of $H$-matrices. This approach allows for the extension of most matrix eigenvalue localization sets to higher-order tensor eigenvalues. Additionally, significant inclusion sets for tensor eigenvalues are introduced, including modified versions of Brauer’s Ovals of Cassini sets, Ostrowski sets, and $S$-type inclusion sets.
5.1 The applications of tensor-generated in $H$-Eigenvalue inclusion sets for high order tensors
[…]
Theorem 11.
[52, Theorem 6] Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$, $n\geq 2$. Then
$\sigma(\mathcal{A})\subseteq\Gamma(\mathcal{A})=\bigcup_{i\in N}\Gamma_{i}(\mathcal{A}),$
where $\Gamma_{i}(\mathcal{A})=\{z\in\mathbb{C}:|z-a_{i\cdots i}|\leq r_{i}(\mathcal{A})\},\quad r_{i}(\mathcal{A})=\sum\limits_{\begin{subarray}{c}i_{2},\cdots,i_{m}\in N,\\ \delta_{ii_{2}\cdots i_{m}}=0\end{subarray}}|a_{ii_{2}\cdots i_{m}}|.$
[…]
Theorem 12.
Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$, $n\geq 2$. Then
${\sigma(\mathcal{A})}\subseteq{\mathcal{F}(\mathcal{A})}=\bigcup_{i\in N}{\mathcal{F}_{i}}(\mathcal{A}),$
where ${\mathcal{F}_{i}}(\mathcal{A})=\{z\in\mathbb{C}:[|z-a_{i\cdots i}|-s_{ii}(\mathcal{A})]\leq\sum\limits_{j=1,j\neq i}^{n}s_{ij}(\mathcal{A})\}$.
Proof.
For any $\lambda\in\sigma(\mathcal{A})$, let $\mathcal{B}=(b_{i_{1}i_{2}\cdots i_{m}})=\lambda\mathcal{I}-\mathcal{A}$. Then
$0\in\sigma(\mathcal{B}),\quad r_{i}(\mathcal{A})=r_{i}(\mathcal{B}),\quad s_{ij}(\mathcal{A})=s_{ij}(\mathcal{B}),\quad\forall i,j\in N.$
We assume $\lambda\notin\mathcal{F}(\mathcal{A})$, which indicate that there is $\lambda\notin\mathcal{F}_{i}(\mathcal{A}),\forall i\in N$. In other words, we obtain
$|\lambda-a_{i\cdots i}|-s_{ii}(\mathcal{A})>\sum\limits_{j=1,j\neq i}^{n}|s_{ij}(\mathcal{A})|,\quad\forall i\in N.$
Equivalently, we have
$|b_{i\cdots i}|-s_{ii}(\mathcal{B})>\sum\limits_{j=1,j\neq i}^{n}|s_{ij}(\mathcal{B})|,\quad\forall i\in N.$
According to equalities (12) and (14), $\mathcal{B}$ is a strictly diagonally dominant tensor. Moreover, there is $0\notin\sigma(\mathcal{B})$, which is in contradiction with the $0\in\sigma(\mathcal{B})$. The proof is completed. ∎
Remark 5.
[…]
Theorem 13 (Brauer’s Ovals of Cassini [60]).
Let $A=(a_{ij})\in\mathbb{C}^{n,n}$ be a $n\times n$ complex matrix, $n\geq 2$, and $\sigma(A)$ be the spectrum of $A$. Then
$\sigma(A)\subseteq\Theta(A)=\bigcup\limits_{\begin{subarray}{c}i,j\in N\\ j\neq i\end{subarray}}\Theta_{ij}(A),$
where $\Theta_{ij}(A)=\{z\in\mathbb{C}:|z-a_{ii}||z-a_{jj}|\leq r_{i}(A)r_{j}(A)\}$, $r_{i}(A)=\sum\limits_{k\neq i}|a_{ik}|$.
However, with the aid of the connections between tensors and tensor-generated matrices, along with the principles of double diagonal dominance and matrix eigenvalue distribution, we are able to derive the following modified Brauer set for higher-order tensor $\mathcal{A}$.
Theorem 14.
Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$. Then
$\sigma(\mathcal{A})\subseteq\mathcal{D}(\mathcal{A})=\bigcup_{\begin{subarray}{c}i,j\in N\\ j\neq i\end{subarray}}\mathcal{D}_{i,j}(\mathcal{A}),$
where
$\mathcal{D}_{i,j}(\mathcal{A})=\{z\in\mathbb{C}:[|z-a_{ii\cdots i}|-s_{ii}(\mathcal{A})][|z-a_{jj\cdots j}|-s_{jj}(\mathcal{A})]\leq P_{i}(A)P_{j}(A)\}.$
Proof.
For any $\lambda\in\sigma(\mathcal{A})$, let $\mathcal{B}=(b_{i_{1}i_{2}\cdots i_{m}})=\lambda\mathcal{I}-\mathcal{A}$. Then
$0\in\sigma(\mathcal{B}),\quad r_{i}(\mathcal{A})=r_{i}(\mathcal{B}),\quad s_{ij}(\mathcal{A})=s_{ij}(\mathcal{B}),\quad\forall i,j\in N.$
We assume $\lambda\notin\mathcal{D}(\mathcal{A})$, which indicate that there is $\lambda\notin\mathcal{D}_{i,j}(\mathcal{A}),\forall i,j\in N$. In other words, we get
$[|\lambda-a_{ii\cdots i}|-s_{ii}(\mathcal{A})][|\lambda-a_{jj\cdots j}|-s_{jj}(\mathcal{A})]>P_{i}(A)P_{j}(A),\quad\forall i,j\in N,$
that is,
$[|b_{ii\cdots i}|-s_{ii}(\mathcal{B})][|b_{jj\cdots j}|-s_{jj}(\mathcal{B})]>P_{i}(B)P_{j}(B),\quad\forall i,j\in N.$
According to equalities (12) and (15), the tensor-generated matrix $B$ of tensor $\mathcal{B}$ is a strictly doubly diagonally dominant matrix, that is, $\mathcal{B}$ is a $H$-tensor. Moreover, there is $0\notin\sigma(\mathcal{B})$, which is in contradiction with the $0\in\sigma(\mathcal{B})$. The proof is completed. ∎
[…]
Theorem 15 (Ostrowski [60]).
Let $A=(a_{ij})\in\mathbb{C}^{n,n}$ be a $n\times n$ complex matrix, $n\geq 2$, $0\leq\gamma\leq 1$, and $\sigma(A)$ be the spectrum of $A$. Then
$\sigma(A)\subseteq\Omega(A)=\bigcup\limits_{i=1}^{n}\Omega_{ij}(A),$
where $\Theta_{ij}(A)=\{z\in\mathbb{C}:|z-a_{ii}|\leq(r_{i}(A))^{\gamma}(c_{i}(A))^{1-\gamma}\}$, $r_{i}(A)=\sum\limits_{j\neq i}|a_{ij}|$, $c_{i}(A)=\sum\limits_{j\neq i}|a_{ji}|$.
While direct generalization of Ostrowski sets to higher-order tensors, akin to matrix eigenvalue Brauer sets, is unfeasible, the modified Ostrowski sets tailored for higher-order tensors $\mathcal{A}\in\mathbb{C}^{[m\times n]}$ can be obtained by leveraging the correlation between the tensor $\mathcal{A}\in\mathbb{C}^{[m\times n]}$ and its corresponding tensor-generated matrix $A\in\mathbb{C}^{n,n}$. This process involves establishing a connection between $\gamma$-diagonally dominant matrices and the distribution of matrix eigenvalues.
Theorem 16.
Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$, $n\geq 2$, $0\leq\gamma\leq 1$. Then
$\sigma(\mathcal{A})\subseteq\mathcal{O}(\mathcal{A})=\bigcup_{i\in N}\mathcal{O}_{i}(\mathcal{A}),$
where
$\mathcal{O}_{i}(\mathcal{A})=\{z\in\mathbb{C}:|z-|a_{i\cdots i}|+s_{ii}(\mathcal{A})|\leq[P_{i}(A)]^{\gamma}[Q_{i}(A)]^{1-\gamma}\}.$
Proof.
For any $\lambda\in\sigma(\mathcal{A})$, let $\mathcal{B}=(b_{i_{1}i_{2}\cdots i_{m}})=\lambda\mathcal{I}-\mathcal{A}$. Then
$0\in\sigma(\mathcal{B}),\quad r_{i}(\mathcal{A})=r_{i}(\mathcal{B}),\quad s_{ij}(\mathcal{A})=s_{ij}(\mathcal{B}),\quad\forall i,j\in N.$
We assume $\lambda\notin\mathcal{O}(\mathcal{A})$, which indicate that there is $\lambda\notin\mathcal{O}_{i}(\mathcal{A}),\forall i\in N$. In other words, we obtain
$[|\lambda-a_{i\cdots i}|-s_{ii}(\mathcal{A})]>[P_{i}(A)]^{\gamma}[Q_{i}(A)]^{1-\gamma},\quad\forall i\in N.$
That is to say
$[|b_{i\cdots i}|-s_{ii}(\mathcal{B})]>[P_{i}(B)]^{\gamma}[Q_{i}(B)]^{1-\gamma},\quad\forall i\in N.$
According to equalities (12) and (16), the tensor-generated matrix $B$ of tensor $\mathcal{B}$ is a strictly $\gamma$-diagonally dominant matrix, that is to say, $\mathcal{B}$ is a $H$-tensor. Moreover, there is $0\notin\sigma(\mathcal{B})$, which is in contradiction with the $0\in\sigma(\mathcal{B})$. The proof is completed. ∎
Lemma 7.
If $p,q>0$ and $0\leq\gamma\leq 1$, then
$p^{\gamma}q^{1-\gamma}\leq\gamma p+(1-\gamma)q.$
The equality holds if and only if $p=q$.
Subsequent to Theorem 16 and Lemma 7, the ensuing Corollary is deduced.
Corollary 4.
Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$, $n\geq 2$, $0\leq\gamma\leq 1$. Then
$\sigma(\mathcal{A})\subseteq\mathcal{W}(\mathcal{A})=\bigcup_{i\in N}\mathcal{W}_{i}(\mathcal{A}),$
where
$\mathcal{W}_{i}(\mathcal{A})=\{z\in\mathbb{C}:[|z-a_{i\cdots i}|-s_{ii}(\mathcal{A})]\leq\gamma P_{i}(A)+(1-\gamma)Q_{i}(A)\}.$
Let set $S$ be any nonempty subset of $N=\{1,2,\cdots,n\}$, $n\geq 2$, where $\overline{S}=N\setminus S$ is the complement of the subset $S$ in $N$. For any given matrix $A=(a_{ij})\in\mathbb{C}^{n,n}$, each row sum $r_{i}(A)$ is divided into $S$ and $\overline{S}$ parts:
$\begin{cases}r_{i}(A)=\sum\limits_{j\in N\setminus\{i\}}|a_{ij}|=r_{i}^{S}(A)+r_{i}^{\overline{S}}(A),\\ r_{i}^{S}(A)=\sum\limits_{j\in S\setminus\{i\}}|a_{ij}|,\quad r_{i}^{\overline{S}}(A)=r_{i}(A)-r_{i}^{S}(A)=\sum\limits_{j\in\overline{S}\setminus\{i\}}|a_{ij}|,i\in N.\end{cases}$
Definition 8.
Let set $S$ be any nonempty subset of $N=\{1,2,\cdots,n\}$, $n\geq 2$, for any given matrix $A=(a_{ij})\in\mathbb{C}^{n,n}$, then the matrix $A$ is strictly $S$-diagonally dominant, if satisfied:
$\begin{cases}|a_{ii}|>r_{i}^{S}(A),\forall i\in S,\\ |a_{ii}|-r_{i}^{S}(A)\big{)}\cdot\big{(}|a_{jj}|-r_{j}^{\overline{S}}(A)\big{)}>r_{i}^{\overline{S}}(A)r_{j}^{S}(A),i\in S,j\in\overline{S}.\end{cases}$
The matrix eigenvalue $S$-type inclusion theorem is a prominent classical theory within the field of matrix eigenvalue inclusion theorems. This theorem directly follows from the matrix eigenvalue inclusion theorem based on $S$-diagonally dominant matrices.
Theorem 17.
[…]
$\Gamma_{i}^{S}(A)=\{z\in\mathbb{C}:|z-a_{ii}|\leq r_{i}^{S}(A)\},\forall i\in S,$
and the sets
$V_{i,j}^{S}(A)=\{z\in\mathbb{C}:\big{(}|z-a_{ii}|-r_{i}^{S}(A)\big{)}\cdot\big{(}|z-a_{jj}|-r_{j}^{\overline{S}}(A)\big{)}\leq r_{i}^{\overline{S}}(A)r_{j}^{S}(A)\},$
($\forall i\in S,\forall j\in\overline{S}$). Then,
$\sigma(A)\subseteq C^{S}(A)=\bigg{(}\bigcup_{i\in S}\Gamma_{i}^{S}(A)\bigg{)}\bigcup\bigg{(}\bigcup_{i\in S,j\in\overline{S}}V_{i,j}^{S}(A)\bigg{)}.$
$\begin{cases}P_{i}(\mathcal{A})=\sum\limits_{j\in N\setminus\{i\}}|a_{ij}|=r_{i}^{S}(\mathcal{A})+r_{i}^{\overline{S}}(\mathcal{A}),\\ r_{i}^{S}(\mathcal{A})=\sum\limits_{j\in S\setminus\{i\}}|s_{ij}|,\quad r_{i}^{\overline{S}}(\mathcal{A})=P_{i}(\mathcal{A})-r_{i}^{S}(\mathcal{A}),i\in N.\end{cases}$
Exploiting the established relationship between tensors and tensor-generated matrices in the previous section, we expand the classical matrix eigenvalue $S$-type inclusion theorem to encompass higher-order tensor $H$-eigenvalues. Consequently, we present a modified $S$-type inclusion theorem specifically tailored for tensor $H$-eigenvalues. This theorem enables us to effectively characterize the inclusion of $H$-eigenvalues within higher-order tensors by leveraging the connection between tensors and their tensor-generated matrices.
Theorem 18.
Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$ be a $m$ order $n$ dimension tensor and $S$ be any nonempty subset of $N=\{1,2,\cdots,n\}$, $n\geq 2$, where $\overline{S}=N\setminus S$ is the complement of the subset $S$ in $N$. Then
$\sigma(\mathcal{A})\subseteq\mathcal{C}^{S}(\mathcal{A})=\bigg{(}\bigcup_{i\in S}\Gamma_{i}^{S}(\mathcal{A})\bigg{)}\bigcup\bigg{(}\bigcup_{i\in S,j\in\overline{S}}\mathcal{C}_{i,j}^{S}(\mathcal{A})\bigg{)}$
where
$\Gamma_{i}^{S}(\mathcal{A})=\{z\in\mathbb{C}:||z-a_{i\cdots i}|-s_{ii}(\mathcal{A})|\leq r_{i}^{S}(\mathcal{A})\},\forall i\in S,$
$\begin{split}\mathcal{C}_{i,j}^{S}(\mathcal{A})=\{z\in\mathbb{C}:&[||z-a_{i\cdots i}|-s_{ii}(\mathcal{A})|-r_{i}^{S}(\mathcal{A})][||z-a_{j\cdots j}|-s_{jj}(\mathcal{A})|-r_{j}^{\overline{S}}(\mathcal{A})]\\ &\leq r_{i}^{\overline{S}}(\mathcal{A})r_{j}^{S}(\mathcal{A})\}.\end{split}$
Proof.
Let $\mathcal{B}=(b_{i_{1}i_{2}\cdots i_{m}})=\lambda\mathcal{I}-\mathcal{A}$, for any $\lambda\in\sigma(\mathcal{A})$, then
$0\in\sigma(\mathcal{B}),\quad r_{i}(\mathcal{A})=r_{i}(\mathcal{B}),\quad s_{ij}(\mathcal{A})=s_{ij}(\mathcal{B}),\quad\forall i,j\in N,$
and
$r_{i}^{S}(\mathcal{A})=r_{i}^{S}(\mathcal{B}),r_{i}^{\overline{S}}(\mathcal{A})=r_{i}^{\overline{S}}(\mathcal{B}),\quad\forall i\in N.$
We assume that $\lambda\notin\mathcal{C}^{S}(\mathcal{A})$, which shows that $\lambda\notin\bigg{(}\bigcup_{i\in S}\Gamma_{i}^{S}(\mathcal{A})\bigg{)}\bigcup\bigg{(}\bigcup_{i\in S,j\in\overline{S}}\mathcal{C}_{i,j}^{S}(\mathcal{A})\bigg{)},\forall i\in N$. In other words, we have
$\begin{cases}||\lambda-a_{i\cdots i}|-s_{ii}(\mathcal{A})|>r_{i}^{S}(\mathcal{A}),\\ [||\lambda-a_{i\cdots i}|-s_{ii}(\mathcal{A})|-r_{i}^{S}(\mathcal{A})][||\lambda-a_{j\cdots j}|-s_{jj}(\mathcal{A})|-r_{j}^{\overline{S}}(\mathcal{A})]>r_{i}^{\overline{S}}(\mathcal{A})r_{j}^{S}(\mathcal{A}),\end{cases}$
equivalently,
$\begin{cases}||b_{i\cdots i}|-s_{ii}(\mathcal{B})|>r_{i}^{S}(\mathcal{B}),\\ [||b_{i\cdots i}|-s_{ii}(\mathcal{B})|-r_{i}^{S}(\mathcal{B})][||b_{j\cdots j}|-s_{jj}(\mathcal{B})|-r_{j}^{\overline{S}}(\mathcal{B})]>r_{i}^{\overline{S}}(\mathcal{B})r_{j}^{S}(\mathcal{B}),\end{cases}$
From Definition 1 and Definition 8, the tensor-generated matrix $B$ of tensor $\mathcal{B}\in\mathbb{C}^{[m\times n]}$ is $S$-strictly diagonally dominant matrix. Hence, $\mathcal{B}$ is a $H$-tensor based on Theorem 1. Furthermore, it is result in $0\notin\sigma(\mathcal{B})$ , which is contradicted with $0\in\sigma(\mathcal{B})$. The proof is completed. ∎
When $S=\{i\}$ for $i\in N$, then Theorem 18 has the following case.
Corollary 5.
Let $\mathcal{A}=(a_{i_{1}i_{2}\cdots i_{m}})\in\mathbb{C}^{[m\times n]}$. Then
$\sigma(\mathcal{A})\subseteq\mathcal{C}^{S^{i}}(\mathcal{A})=\bigcup_{i\in S,j\in\overline{S}}\mathcal{C}_{i,j}^{S^{i}}(\mathcal{A})$
where
$\mathcal{C}_{i,j}^{S^{i}}(\mathcal{A})=\{z\in\mathbb{C}:||z-a_{i\cdots i}|-s_{ii}(\mathcal{A})|\big{(}||z-a_{j\cdots j}|-s_{jj}(\mathcal{A})|-P_{j}(\mathcal{A})+s_{ji}(\mathcal{A})\big{)}\leq P_{i}(\mathcal{A})s_{ji}(\mathcal{A})\}.$
[…]
Theorem 4 reveals the consistent relationship between the irreducibility of the tensor-generated matrix and the weak irreducibility of the original tensor. Consequently, when dealing with irreducible tensor-generated matrices, our method is also suitable for Brualdi-type inclusion sets [70, 60]. Although we will omit the derivation process here, similar principles are applied.
To summarize, we transform the problem of tensor eigenvalue localization into matrix eigenvalue localization. This allows us to leverage the mature theory of matrix eigenvalue localization and apply it to the investigation of tensor eigenvalue localization.
5.2 Numerical examples
Finally, we provide numerical examples to demonstrate the validity and effectiveness of the tensor $H$-eigenvalue inclusion theorem derived in the above conclusions.
We analyze Example 1.1 from [67], demonstrating that Brauer’s Ovals of Cassini sets, typically linked with matrix eigenvalues, do not directly apply to an inclusion theorem for higher-order tensor $H$-eigenvalues.
Example 1.
Let $\mathcal{A}\in\mathbb{R}^{[4\times 2]}$, where the non-zero elements are as follows:
$a_{1111}=7,a_{1112}=a_{1121}=a_{1211}=a_{2111}=-2;$
$a_{2222}=6,a_{2221}=a_{2212}=a_{2122}=a_{1222}=-1;$
otherwise $a_{i_{1}i_{2}i_{3}i_{4}}=0$.
For a 4 order 2 dimension tensor, we can relate the elements of this tensor to a $2\times 2$ tensor-generated matrix (6) due to the connection between the tensor and the tensor-generated matrix. Hence, we obtain:
$A=\left(\begin{array}[]{cc}|7|-|4|&3\\ 3&|6|-|2|\\ \end{array}\right).$
According to Theorem 14, there are
$\sigma(\mathcal{A})\subseteq\mathcal{\bar{O}}(\mathcal{A})=\{z\in\mathbb{C}:0.4586\leq z\leq 12.8541\}.$
In fact, the eigenvalues of tensor $\mathcal{A}=(a_{i_{1}i_{2}i_{3}i_{4}})$ are $\sigma(\mathcal{A})=\{12.7389,0.4725\}.$ This demonstrates the accuracy of the adjusted ovality theorem for higher-order tensor eigenvalues.
Example 2.
Consider $\mathcal{A}\in\mathbb{R}^{[4\times 4]}$ as a real tensor of order 4 and dimension 4 with specific element assignments:
$a_{1111}=10,\quad a_{2222}=8,\quad a_{3333}=7,\quad a_{4444}=5,\quad a_{1333}=a_{1444}=1,$
$a_{1211}=a_{1113}=a_{1141}=1,\quad a_{1332}=a_{1442}=a_{1232}=a_{1234}=a_{1321}=a_{1214}=1,\quad a_{2333}=a_{2444}=1,$
$\quad a_{2112}=a_{2234}=a_{2113}=a_{2343}=a_{2123}=1,\quad a_{3222}=1,\quad a_{3111}=1,\quad a_{3121}=a_{3434}=a_{3123}=1,$
$a_{4222}=1,\quad a_{4111}=1,\quad a_{4121}=a_{4334}=1,\quad\text{and other entries are zero.}$.
The elements of $\mathcal{A}$ in $\mathbb{R}^{[4\times 4]}$ can be associated with a $4\times 4$ tensor-generated matrix (6), resulting in:
$A=\left(\begin{array}[]{cccc}10-\frac{8}{3}&\frac{8}{3}&3&\frac{8}{3}\\ \frac{5}{3}&8-1&\frac{8}{3}&\frac{5}{3}\\ 2&\frac{5}{3}&7-\frac{2}{3}&\frac{2}{3}\\ \frac{5}{3}&\frac{4}{3}&\frac{2}{3}&5-\frac{1}{3}\\ \end{array}\right).$
The numerical results from Theorem 12, Theorem 14, Corollary 4, Theorem 18, and Corollary 5, as well as findings from [52, 67, 71, 72], are presented in Table 1. This illustrates that our eigenvalue localization set, encompassing Theorem 12, Theorem 14, Corollary 4, Theorem 18, and Corollary 5, effectively captures all eigenvalues. Additionally, results from our methods in the mentioned examples reveal the enhanced precision of modified Brauer-type, Ostrowski, and $S$-type inclusion sets for tensor eigenvalues from the tensor-generated matrix compared to existing sets. Specifically, all the $H$-eigenvalues of the tensor $\mathcal{A}$ can be computed as follows: $\sigma(\mathcal{A})=\{4.4858,7.3107,9.7718,15.2641\}.$
Table 1: Lower and upper bounds of $\sigma(\mathcal{A})$.
  Method   Lower bounds   Upper bounds
Theorem 6 in [52] -1 21
Theorem 12 -1 21
Theorem 2.2 in [67] -0.7016 20.3739
Theorem 4 in [71] -0.4641 19.7823
Theorem 11 in [72] -0.3723 19.6847
Theorem 14 0.0936 18.1382
Theorem 16, $\gamma=0.5$ 0.3849 19.3333
Theorem 16, $\gamma=0.04$0.2717 18.0961
Corollary 4,$\gamma=0.5$ 0.3333 19.5000
Corollary $\gamma=0.04$0.2800 18.1200
Theorem 18,$S=\{1,2\}$ 0.5811 17.5803
Corollary 5 -0.4741 19.8130
Conclusions
Multipartite quantum scenarios serve as crucial and complex resources within the domain of quantum information science. Tensors establish a robust framework for the representation of multipartite quantum systems. This study introduces the novel concept of tensor-generated matrices, which define the relationships between an $m$-th order $n$-dimensional tensor and an $n$-dimensional square matrix. By establishing these connections, we illustrate that designating the tensor-generated matrix as an $H$-matrix implies the classification of the original tensor as an $H$-tensor. Additionally, we explore various analogous properties shared by both the original tensor and the tensor-generated matrix, encompassing attributes like weak irreducibility, weakly chained diagonal dominance, and (strong) symmetry. These discoveries offer a methodology to convert intricate tensor issues into matrix formulations within specific contexts, a particularly pertinent approach given the inherent NP-hard complexity of most tensor problems.
Subsequently, we delve into the utilization of tensor-generated matrices to analyze the classicality of spin states. Capitalizing on the tensor representation, we introduce classicality criteria tailored for (strongly) symmetric spin-$j$ states, potentially introducing novel perspectives into the examination of multipartite quantum resources. Furthermore, we extend classical matrix eigenvalue inclusion sets to higher-order tensor $H$-eigenvalues, a challenge often encountered with higher-order tensors. Consequently, we propose representative tensor $H$-eigenvalue inclusion sets, such as modified versions of Brauer’s Ovals of Cassini sets, Ostrowski sets, and $S$-type inclusion sets.
Data availibility
All data generated or analysed during this study are included in this published article.
Declarations
Conflict of interest The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.